### v1

# Surg-3M: A Dataset and Foundation Model for Perception in Surgical Settings

Abstract
Advancements in computer-assisted surgical procedures heavily rely on accurate visual data interpretation from camera systems used during surgeries. Traditional open-access datasets focusing on surgical procedures are often limited by their small size, typically consisting of fewer than 100 videos with less than 100K images. To address these constraints, a new dataset called Surg-3M has been compiled using a novel aggregation pipeline that collects high-resolution videos from online sources. Featuring an extensive collection of over 4K surgical videos and more than 3 million high-quality images from multiple procedure types, Surg-3M offers a comprehensive resource surpassing existing alternatives in size and scope, including two novel tasks. To demonstrate the effectiveness of this dataset, we present SurgFM, a self-supervised foundation model pretrained on Surg-3M that achieves impressive results in downstream tasks such as surgical phase recognition, action recognition, and tool presence detection. Combining key components from ConvNeXt, DINO, and an innovative augmented distillation method, SurgFM exhibits exceptional performance compared to specialist architectures across various benchmarks. Our experimental results show that SurgFM outperforms state-of-the-art models in multiple downstream tasks, including significant gains in surgical phase recognition (+8.9pp, +4.7pp, and +3.9pp of Jaccard in AutoLaparo, M2CAI16, and Cholec80), action recognition (+3.1pp of mAP in CholecT50) and tool presence detection (+4.6pp of mAP in Cholec80). Moreover, even when using only half of the data, SurgFM outperforms state-of-the-art models in AutoLaparo and achieves state-of-the-art performance in Cholec80. Both Surg-3M and SurgFM have significant potential to accelerate progress towards developing autonomous robotic surgery systems. Our dataset, models, and code are publicly available at https://github.com/visurg-ai/surg-3m.
1 Introduction
Figure 1: Performance comparison between our foundation model, SurgFM, and the state-of-the-art (SotA) models. Our evaluation focuses on three surgical downstream tasks and six datasets. SurgFM results are shown in bold. Axis labels are written in regular font. For a detailed breakdown of the results, see Tables 3, 4, and 5.
Figure 2: Diversity and procedure prevalence in Surg-3M. Representative samples from various procedures, demonstrating the diverse range of cases in our curated dataset (left, right). Distribution of surgical frames by procedure type (center).
Autonomous surgery has long been a promising area of research, with the potential to revolutionize the way surgical procedures are performed [25]. To achieve this goal, significant advancements in perception and computer vision are necessary, allowing robotic systems to accurately interpret and respond to the complexities of surgical environments [38]. However, training models for surgical perception is far from trivial due to the limited availability and high quality requirements of annotated data [59, 39, 45]. The complexities of medical data and stringent regulations governing patient privacy create significant hurdles for collecting and annotating relevant information [19, 2]. The cumulative effect of all these challenges has resulted in significantly limited computer vision datasets for surgical applications, with many containing fewer than 10K images [50]. These datasets, which may suffice for validation on specific computer vision tasks and surgical procedures, fall short in providing an adequate foundation for training models that can generalize effectively across diverse surgical settings and datasets.
Recent advances in self-supervised learning have opened up new research avenues, pretrained foundation models can significantly expedite progress, potentially reducing the need for annotated data by utilizing large-scale unlabeled datasets [49]. Despite this progress, ongoing data collection challenges persist, motivating the exploration of two key aspects: the sufficiency of publicly available surgical videos as a foundation for building a comprehensive dataset, and the potential of self-supervised models trained on this dataset to achieve state-of-the-art performance in surgical downstream tasks. By addressing these aspects, we aim to provide valuable insights into the feasibility and efficacy of using public domain data for training machine learning models in the field of surgical video analysis.
Figure 3: Data curation pipeline. (a) Surgical videos are collected from YouTube. (b) The videos are summarized as storyboards (trivial to annotate and classify). We train a storyboard classifier to identify videos with rich surgical content, and manually verify the selected videos. (c) During the video selection and trimming phase, a trained frame classifier categorizes each frame as either surgical or non-surgical, allowing us to trim non-surgical frame segments at the beginning and end of each video. We further filter videos by requiring at least 90% of their frames to be surgical frames. (d) In the video preprocessing stage, we obliterate non-surgical frames. (e) During the video annotation phase, we utilize the video title as a primary cue to determine the procedure and surgery type. In cases where the video titles do not explicitly include any of our procedure name keywords (Fig. 2), we employ ChatGPT to match the titles with surgical procedure types. The set of final videos, and their corresponding labels, are manually quality controlled.
Contributions. We propose a novel data curation pipeline (Fig. 3) for collecting videos of selected surgical procedures, capturing exclusively surgical footage while discarding non-surgical content. Building upon this data curation pipeline, we present Surg-3M, a dataset for computer vision applications in surgery. Our Surg-3M dataset is large and diverse, containing 4K surgical videos and 3M images (Table 1). To the best of our knowledge, Surg-3M is the largest open-access surgical video dataset to date, 18×18$\times$18 × larger than the largest published surgical dataset [54]. To model the features of surgical video frames, we provide SurgFM, a pretrained model based on the ConvNeXt architecture [36] and DINO [8]. Additionally, we propose a novel augmented distillation method (Fig. 4) designed to enhance the local-to-global self-distillation mechanism described in  [8]. SurgFM outperformed the current state-of-the-art models (Fig. 1) on Cholec80 [54], M2CAI16 [54, 53], AutoLaparo [57], and CholecT50 [44]. Additionally, it helps to significantly reduce the annotation effort (Fig. 5). Building on our Surg-3M annotations, we introduce two new supervised learning applications: multi-label surgical procedure classification and surgery type classification. Leveraging our image foundation model, SurgFM, we propose SurgFM-Vid, a video classifier that achieves top performance on these new tasks, serving as a baseline for future research.
Table 1: Comparison of Surg-3M’s size to other existing public surgical datasets. The number of frames is obtained by sampling videos at one frame per second.
Dataset Year # Procedure types # Videos # Frames
M2CAI16 [54] $2016$ 1 41 $95$K
Cholec80 [54] $2016$ 1 80 $183$K
CholecT50 [44] $2022$ 1 50 $101$K
AutoLaparo [57] $2022$ 1 21 $83$K
SAR-RARP50 [48] $2023$ 1 50 $163$K
GraSP [4] $2024$ $1$ $13$ $117$K
Surg-3M (ours) $\mathbf{2025}$ $\mathbf{35}$ $\mathbf{4194}$ $\mathbf{3}$M
2 Related Work
Surgical datasets. Previous works have released open-access datasets aimed at addressing specific surgical downstream tasks and procedures (Table 1). Despite their valuable contributions, these open-access datasets suffer from a narrow scope of surgical procedures and limited dataset sizes. Besides these open-access datasets, there are several private surgical datasets, including SVL [63], with 1.4K videos, and SAGES CVS [47], with 1K videos. However, the restricted accessibility of these datasets constrains their utility and hinders their potential for more widespread use. To overcome these limitations, our work introduces Surg-3M. Compared with Cholec80 [54], the largest open-access surgical dataset, Surg-3M provides 18$\times$ more surgical images and 35$\times$ the surgical procedure types.
Self-supervised learning. Self-supervised learning (SSL) has gained significant attention due to its ability to reduce reliance on large-scale annotated datasets, a benefit that is particularly relevant for surgical applications. Recently, Ramesh et al. [49] investigated the effectiveness of different SSL approaches for surgical applications, including MoCov2 [11], SwAV [7], SimCLR [10], and DINO [8]. Their experiments showed that DINO consistently outperformed other SSL methods on both surgical phase recognition and surgical tool presence detection tasks, highlighting its potential in surgical applications.
As part of a broader effort to advance self-supervised surgical models, Hirsch et al. [24] built a private surgical dataset and trained Masked Siamese Networks (MSN) [3] on this dataset, achieving near-state-of-the-art results. While the research presented in this study shows promising results, its broader applicability and reproducibility are somewhat constrained by the lack of public availability of the underlying dataset, as well as a relatively narrow evaluation scope that only considers performance on two specific task-dataset pairs: surgical phase recognition on Cholec80 [54], and polyp characterization on PolypSet [32]. Similarly, in Endo-FM [58], the authors proposed to train a student-teacher foundation model on a composite dataset that combined private and public data sources. Although showing encouraging results, the approach is also of limited accessibility. More recently, Batic et al. [5] created a dataset of 70K images merging nine public datasets. They trained a model based on masked autoencoders (MAE)  [23], but showed limited performance, likely due to their reduced dataset size.
[…]
Method Phase recognition (Cholec80) Tool detection (Cholec80) Action recognition (CholecT50)
Accuracy (%) F1-score (%) mAP (%) F1-score (%) mAP (%) F1-score (%)
DINO (frozen) [8] $69.8\pm 1.5$ $62.5\pm 1.5$ $67.8\pm 3.1$ $59.3\pm 4.0$ $36.1\pm 1.4$ $28.9\pm 2.7$
DINOv2 (frozen) [46] $67.0\pm 1.7$ $60.3\pm 1.8$ $56.1\pm 3.4$ $41.2\pm 2.6$ $27.2\pm 2.2$ $17.2\pm 0.8$
MAE (frozen) [23] $55.2\pm 1.4$ $49.6\pm 1.2$ $40.4\pm 1.4$ $22.9\pm 0.9$ 20.7 $\pm$ 1.3 13.4 ±$\pm 1.3$
[…]
3 Proposed Surg-3M Dataset
To develop a robust surgical foundation model, it is essential to create a comprehensive, large-scale dataset of diverse surgical images that covers the whole spectrum of minimally invasive surgical procedures executed within the operating room. To fulfill this need we propose Surg-3M, a dataset that contains 4K videos and 3M images corresponding to 35 different surgical procedures, including videos from robotic-assisted surgeries and non-robotic traditional laparoscopies (more details on these surgical techniques are provided in Sec. A of the supplementary material).
Ground truth and proposed new tasks. Each of the 4K surgical videos is annotated for two tasks: 1) multi-label procedure type classification, where the objective is to identify the procedure[s] performed in an intervention’s video from among 35 possible options (Fig. 2); and 2) binary surgery type classification, where the goal is to indicate whether a surgery is conducted robotically or non-robotically. The rationale for these annotations is two-fold. Firstly, it allows us to validate the comprehensiveness of our dataset by ensuring that it encompasses a broad spectrum of surgical procedures. Secondly, it facilitates the development of new methods that use this information as a key input (e.g., our SurgFM foundation model, video classification/retrieval, video similarity measurement, procedure-specific estimation of surgical remaining time).
Scenes in surgical videos unfold at a relatively slow pace, as surgeons must ensure patient safety. Consequently, our 3M image dataset was compiled by sampling our dataset of 4194 curated videos at 1 fps, which is standard practice in the field of surgical computer vision [54, 57, 44, 48, 53]. In the following sections we detail how we collected, curated, and annotated these videos. An overview of the curation pipeline is shown in Fig. 3.
3.1 Dataset Curation
Video collection. Inspired by works in other fields [15, 9, 61, 1, 40], we collected our videos from YouTube. For each of the $35$ procedures in Fig. 2, we searched for videos of their robotic and non-robotic variants using the strings “robotic $<$procedure type$>$” and “laparoscopic $<$procedure type$>$”, respectively. We run the video collection process until we had $500$ different videos collected for each procedure leading to a total of $\approx$18K raw videos.
Video classification. The raw videos collected based on search keywords contained a diverse range of content, including patient testimonials and conference presentations. To avoid including non-surgical videos in our dataset, we classified the videos as surgical and non-surgical (Fig.  3b). This pipeline consisted of three stages. First, a dataset of 2160 surgical and 1910 non-surgical video storyboards (Fig. 3b) was annotated. Second, a binary video storyboard classifier was trained on this dataset and run over all the raw videos. Finally, manual verification was performed on the videos classified as surgical.
Video selection and trimming. A video classified as surgical may still contain non-surgical frames (e.g., presentation slides). Therefore, we trained a surgical frame classifier and used it to 1) trim the non-surgical content present at the beginning and end of videos, and 2) discard videos containing more than 10% of the frames estimated as non-surgical after trimming (Fig. 3c).
Video preprocessing. We use our previously trained frame classifier to detect and obliterate those frames that are estimated to be non-surgical within the trimmed videos. After this process, we manually quality control the videos to verify that their frames are surgical.
Video annotation. We cross-reference the video titles with the predefined list of procedures. When exact title matches are not found, we leverage the capabilities of the ChatGPTv4 API to perform a more nuanced analysis, incorporating a customized prompt as specified in Sec. B of the supplementary material. All annotations are manually reviewed and validated to ensure the accuracy of the labels assigned.
4 Proposed SurgFM Foundation Model
In this section, we describe how we build SurgFM, our foundation model trained on Surg-3M, and motivate its design choices. This section includes two ablation studies to identify the optimal self-distillation configuration and the most suitable backbone architecture for surgical videos; a description of our image preprocessing pipeline, and our proposed augmented distillation mechanism with its corresponding ablation study. The results of the ablation experiments are presented in Table 2.
Self-supervised methods. We examine the impact of the different self-supervision mechanisms (Table 2). We evaluate DINO [8], DINOv2 [46], and MAE [23]. For these experiments, we trained the models in their standard configurations, a ViT-B backbone [14] pretrained on ImageNet-1k [13] with frozen weights and a linear classification head. Stemming from these results, which align with those in the literature [49], DINO consistently outperforms DINOv2 and MAE on all downstream tasks, with surgical tool detection and action recognition showing a large margin. Therefore, we adopt DINO’s self-distillation [8] as our pretext task for self-supervision.
Backbone models. As noted in [56], ViTs [14] face challenges to capture fine details, which is essential to solving tasks such as tool presence detection and surgical action recognition. Given this, we opted to evaluate the performance of ViT-L [14] against a modern convolutional architecture, ConvNeXt-L [36]. Both were pretrained on Surg-3M and fine-tuned on Cholec80 [54] and CholecT50 [42], for surgical phase recognition, tool presence detection, and surgical action recognition. The results are shown in Table 2. ConvNeXt-L outperformed ViT-L, suggesting that convolutional networks are indeed better equipped to capture those fine image details that are relevant to solve surgical tasks. Therefore, we use ConvNeXt-L as the backbone of our SurgFM model.
Image preprocessing pipeline. Surgical videos often exhibit superimposed manufacturer-specific user interface (UI) information generated by the software of the surgical stack. While certain researchers might derive value from analyzing these UI elements (e.g., some interfaces provide contextual information regarding the surgical instruments utilized in the video), a foundational model for surgery should prioritize learning correlations between tissue appearance and tool-tissue interactions. This focus is essential to avoid acquiring spurious correlations with manufacturer-specific UI elements. Therefore, we manually annotated 2719 surgical frames with 4584 non-surgical bounding-box instances and trained a YOLOv8 Nano model [31] to detect and crop out UI content (see Fig. S2 in supplementary material).
[…]
Method Type Cholec80 AutoLaparo M2CAI16 GraSP
Accuracy (%) Precision (%) Recall (%) Jaccard (%) F1-score (%) Accuracy (%) Precision (%) Recall (%) Jaccard (%) Accuracy (%) Precision (%) Recall (%) Jaccard (%) mAP (%)
ConvNeXt-L [36] G $91.4\pm 6.7$ $92.3\pm 3.8$ $90.5\pm 7.7$ $82.1\pm 7.1$ $88.6\pm 6.8$ $82.3\pm 9.7$ $78.3\pm 21.9$ $69.9\pm 31.9$ $60.5\pm 27.2$ $88.3\pm 10.0$ $89.4\pm 4.7$ $89.5\pm 3.3$ $78.4\pm 6.2$ $73.4$
EndoNet [54] S $81.7\pm 4.2$ $73.7\pm 16.1$ $79.6\pm 7.9$ $76.5$
MTRCNet-CL [27] $89.2\pm 7.6$ $86.9\pm 4.3$ $88.0\pm 6.9$ $87.4$
PhaseNet [52] $78.8\pm 4.7$ $71.3\pm 15.6$ $76.6\pm 16.6$ $73.8$ $79.5\pm 12.1$ $64.1\pm 10.3$
SV-RCNet [26] $85.3\pm 7.3$ $80.7\pm 7.0$ $83.5\pm 7.5$ $82.1$ 75.6 $64.0$ $59.7$ $47.2$ 81.7 $\pm$ 8.1 81.0 $\pm$ 8.3 81.6 $\pm$ 7.2 65.4 $\pm$ 8.9 -
OHFM [62] $87.3\pm 5.7$ $67.0\pm 13.3$ $85.2\pm 7.5$ $68.8\pm 10.5$
TeCNO [12] $88.6\pm 7.8$ $86.5\pm 7.0$ $87.6\pm 6.7$ $75.1\pm 6.9$ $77.3$ $66.9$ $64.6$ $50.7$ $86.1\pm 10.0$ $85.7\pm 7.7$ $88.9\pm 4.5$ $74.4\pm 7.2$
TMRNet [29] $90.1\pm 7.6$ $90.3\pm 3.3$ $89.5\pm 5.0$ $79.1\pm 5.7$ $78.2$ $66.0$ $61.5$ $49.6$
AVT [20] - - - - - $77.8\pm 9.4$ $68.0$ $62.2$ $50.7$
Trans-SVNet [30] $90.3\pm 7.1$ $90.7\pm 5.0$ $88.8\pm 7.4$ $79.3\pm 6.6$ $78.3$ $64.2$ $62.1$ $50.7$ $87.2\pm 9.3$ $88.0\pm 6.7$ $87.5\pm 5.5$ $74.7\pm 7.7$
LoViT [35] $92.4\pm 6.3$ $89.9\pm 6.1$ $90.6\pm 4.4$ $81.2\pm 9.1$ $81.4\pm 7.6$ $85.1$ $65.9$ $55.9$
SlowFast± [17] - - - - - - - - - - - - - $70.1$
TAPIR± [55] - - - - - - - - - - - - - $74.6$
TAPIS± [4] - - - - - - - - - - - - - $\mathbf{76.1}$
Ramesh et al. [49] F - - - - $81.6$
SSL4EVA [24] - - - - $86.9$
EndoViT [5] 89.4 - - - - - - - - - - - - -
SurgFM (ours) $87.0\pm 10.3$ $87.3\pm 7.5$ $87.7\pm 4.3$ $75.7\pm 9.2$ $83.1\pm 11.0$ $81.7\pm 10.9$ $\mathbf{86.3\pm 8.9}$ $69.8\pm 30.9$ $60.2\pm 26.1$ $82.6\pm 10.7$ $85.5\pm 9.9$ $87.3\pm 5.9$ $73.2\pm 9.4$ $66.8$
SurgFM (ours) $\mathbf{92.7\pm 8.3}$ $\mathbf{93.0\pm 4.5}$ $\mathbf{93.1\pm 4.7}$ $\mathbf{85.1\pm 7.9}$ $\mathbf{90.6\pm 6.7}$ $\mathbf{85.2\pm 9.9}$ $82.9\pm 17.3$ $\mathbf{73.2\pm 32.0}$ $\mathbf{64.8\pm 27.3}$ $\mathbf{89.9\pm 7.1}$ $\mathbf{89.6\pm 6.5}$ $\mathbf{89.9\pm 6.9}$ $\mathbf{79.4\pm 9.9}$ $75.9$
Proposed augmented distillation method. Inspired by DINO [8], we train a student network $f_{\boldsymbol{\theta_{s}}}$ to solve for the mapping $\sigma(f_{\boldsymbol{\theta_{t}}}(\boldsymbol{x}))=\sigma(f_{\boldsymbol{\theta_{s}}}(\boldsymbol{x}))$ where $\boldsymbol{x}\in\mathbb{R}^{H\times W\times 3}$ is an input image, $\sigma$ is the softmax function, $f_{\boldsymbol{\theta_{t}}}(\boldsymbol{x})\in\mathbb{R}^{C}$ is a target vector of dimension $C=2^{16}$ estimated from a teacher network $f_{\boldsymbol{\theta_{t}}}$, and $\boldsymbol{\theta_{s}}$ and $\boldsymbol{\theta_{t}}$ are the parameters of the student and teacher models, respectively. We use the outputs of both networks to build two categorical distributions, $P_{s}(z|\sigma(f_{\boldsymbol{\theta_{s}}}(\boldsymbol{x})))=\sigma_{z}(f_{\boldsymbol{\theta_{s}}}(\boldsymbol{x}))$ and $P_{t}(z|\sigma(f_{\boldsymbol{\theta_{t}}}(\boldsymbol{x})))=\sigma_{z}(f_{\boldsymbol{\theta_{t}}}(\boldsymbol{x}))$, where $\sigma_{z}$ denotes the $z^{th}$ element of the softmax output, and $z\in\{1,\dots,C\}$. In training, we minimize the negative log-likelihood of $P_{s}(z|\sigma(f_{\boldsymbol{\theta_{s}}}(\boldsymbol{x})))$ following
$\small\min_{\boldsymbol{\theta_{s}}}\sum_{i=1}^{I}$ $\sum_{\begin{subarray}{c}\boldsymbol{u}\in U_{i}\\ \boldsymbol{v}\in V_{i}\cup W_{i}\\ \boldsymbol{u}\neq\boldsymbol{v}\end{subarray}}\Big{[}d(\boldsymbol{u},\boldsymbol{v})-n(\boldsymbol{u},\boldsymbol{v})\Big{]}$ (1)
where $d(\boldsymbol{u},\boldsymbol{v})=\log\sum_{k^{\prime}=0}^{C-1}\exp\left(\sigma_{k^{\prime}}(f_{\boldsymbol{\theta_{s}}}(\boldsymbol{v}))\right)$, $n(\boldsymbol{u},\boldsymbol{v})=\sigma_{\arg\max_{z}P_{t}(z|\sigma(f_{\boldsymbol{\theta_{t}}}(\boldsymbol{u})))}(f_{\boldsymbol{\theta_{s}}}(\boldsymbol{v}))$, $I$ is the number of images in the dataset, $U_{i}$ contains two global augmentations of $\boldsymbol{x_{i}}$, and $V_{i}$ contains two global augmentations and four augmented crops of $\boldsymbol{x_{i}}$. A foundation model for surgery should learn invariance to 1) minor movements of instruments and tissues observed between adjacent video frames, and 2) minimal appearance changes across patients undergoing the same surgical procedure (e.g., organ color variations). Diverging from DINO  [8], and relying on the annotations available in Surg-3M, we introduce $W_{i}$ to provide a supervisory signal that encourages the model to learn these invariances. $W_{i}$ contains two images, and four augmented crops of these two images. The procedure for selecting these two images is depicted in Fig. 4. We use the $k$-nearest neighbors (in embedded space) of the input image from other videos of the same procedure type to learn invariance to appearance changes across patients, and the prior and subsequent video frames of the input image to learn invariance to the motion occurring between consecutive frames.
We evaluate the performance of models trained with and without our augmented distillation (see Table 2) on surgical phase recognition (Cholec80), tool presence detection (Cholec80), and surgical action recognition (CholecT50). With augmented distillation, F1-score improved by $>1pp$ in both tool detection and action recognition tasks.
Following these ablation results, we incorporated augmented distillation in our final SurgFM model.
Figure 4: Proposed augmented distillation method. To encourage invariance in SurgFM to minor surgical motion and subtle appearance changes across surgeries, we introduce $W_{i}$ (Eq. 1). The key component of $W_{i}$ is a pair of images (c) that are randomly selected from an augmentation pool (b). To populate the augmentation pool, which has capacity for four images, we first retrieve the nearest neighbors(1) of the input image from other videos of the same procedure type, but only include them in the pool if the cosine distance to the input image is smaller than 3×3$\times$3 × the distance between the input image and its preceding frame in the video that it was extracted from. When not enough suitable neighbors are found, we supplement the pool with adjacent video frames(2)(3).
Proposed video classification model. Considering that each surgical procedure is localized to a specific region of the body, we posit that a characteristic scene can be associated with each procedure. Based on this assumption, and inspired by active learning [21], we propose SurgFM-Vid, a video classification method that employs SurgFM embeddings for video classification by aggregating them following
$\small\boldsymbol{v_{e}}=\sum_{j=0}^{J-1}\omega_{j}\boldsymbol{\phi}_{j}$ (2)
where $\boldsymbol{v_{e}}$ is a video embedding, $J$ is the number of frames in the video, $\omega_{j}=\frac{t(\boldsymbol{\phi}_{j})}{\epsilon+\sum_{m=0}^{J-1}t(\boldsymbol{\phi}_{m})}$, $\boldsymbol{\phi}_{j}$ is the embedding of frame $j$, $t(\boldsymbol{\phi})=\left(\frac{1}{K}\sum_{\boldsymbol{\eta}\in K-NN(\boldsymbol{\phi})}d(\boldsymbol{\phi},\boldsymbol{\eta})\right)^{-1}$ is the typicality [21] of a frame, $K-NN(\boldsymbol{\phi})$ with $K=20$ are the nearest neighbors of $\boldsymbol{\phi}$, $d$ is the cosine distance, and $\epsilon=10^{-8}$. Then, each $\boldsymbol{v_{e}}$ is classified with a single-layer MLP.
5 Evaluation
5.1 Downstream Tasks, Datasets and Metrics
We evaluate SurgFM on three surgical downstream tasks: surgical phase recognition [54, 53, 57, 4], surgical tool presence detection [54, 4], and surgical action recognition [48, 4]. To do so, we use the most widely adopted datasets, data splits, and evaluation metrics.
Surgical phase recognition. We use M2CAI16 [54], Cholec80 [54], AutoLaparo [57], and GraSP [4]. For this task, we report accuracy, precision, recall, Jaccard, F1-score, and mean Average Precision (mAP). The accuracy for a video is computed by dividing the number of frames whose class has been correctly predicted by the total number of frames. The overall accuracy for a dataset is defined as the mean accuracy value over all videos, henceforth referred to as video-level accuracy [28, 29, 35, 27]. Precision, recall, and Jaccard are computed differently than accuracy [28, 29, 35, 27]. These three phase-level metrics are computed for each class and video, then averaged: first within videos, and secondly across classes. The F1-score, computed for Cholec80, is also computed for each class and video, but averaged first over classes and then over videos, and referred to as video-level F1-score [49, 24]. For GraSP, the frame-level mAP was calculated by averaging the mAP scores across all frames in the test set [4].
Surgical tool presence detection. Cholec80 and GraSP are utilized in the surgical phase recognition task. We report mAP across all frames in the testing dataset [49, 4].
Surgical action recognition. For this task we use CholecT50 [44] and SAR-RARP50 [48]. For CholecT50, we report the video-level mAP, which is computed by first calculating the average mAP for each class across all videos and then taking the overall mean of these class averages [44]. For SAR-RARP50 [48], we report video-level accuracy and video-level F1@10 [48].
5.2 Downstream Task Evaluation Protocols
To capture temporal correlations in the surgical phase recognition downstream task, we couple our SurgFM backbone with a TCN head [12]. This TCN head takes both current and previous frames into account, but excludes future frames. For a fair comparison with previous methods [26, 54, 30, 12, 27, 35], we run our evaluation in Cholec80 and M2CAI16 with the standard 10-second relaxed boundary between some of the surgical phases [54]. Also to be able to compare with previous literature [57, 35, 4], we do not use a relaxed boundary for AutoLaparo and GraSP.
For surgical tool presence detection, we use a linear head for all datasets.
For surgical action recognition we use a different head for each dataset. To evaluate on CholecT50, we utilize a linear head. The predictions are done per-frame, no temporal information is used, as in the approaches that we benchmark against [44, 43]. For SAR-RARP50, we evaluate our model in two different modes: online, where only previous frames are available (TCN head [12]), and offline ($\pm$ subscript in Table 5), which incorporates information from both previous and future frames (MS-TCN head [33]), to ensure a fair comparison with all competing methods in [48].
In our evaluation, we compare three types of models: generalist (G), pretrained on ImageNet and fine-tuned for each downstream task; specialist (S), designed for specific downstream tasks; and foundational (F), pretrained in a self-supervised manner and fine-tuned on each downstream task.
Table 4: Performance evaluation on surgical tool presence detection. We evaluate the performance on two datasets: Cholec80 [54], and GraSP [4]. The metrics are computed as described in Sec. 5.1. The evaluation protocol and the meanings of G, S, and F in the type column are described in Sec. 5.2. The superscript denotes that the backbone is frozen during fine-tuning.
Method Type Cholec80 GraSP
mAP (%) mAP (%)
ConvNeXt-L [36] G 69.3 $68.1$
ConvNeXt-L [36] $93.3$ $89.5$
EndoNet [54] S $81.0$
MTRCNet-CL [27] $89.1$
Rendezvous [44] - $80.4$
TAPIS [4] - $\mathbf{95.1}$
Ramesh et al. [49] F $93.6$
SurgFM (ours) $76.1$ $76.4$
SurgFM (ours) $\mathbf{93.7}$ $94.4$
6 Results and Discussion
6.1 Ablation Studies
SurgFM design choices. We performed incremental ablation studies that test the choice of backbone (ViT vs ConvNeXt), self-supervision method (DINO vs DINOv2 vs MAE), and our proposed augmented distillation (Sec. 4). The results are shown in Table 2. DINO outperformed DINOv2 and MAE in all the downstream tasks and metrics evaluated, with improvements $>$8 percentage points (pp) in mAP in both tool detection and action recognition. ConvNeXt-L achieved comparable results to ViT-L in phase recognition, and improved $>$2pp (F1-score) in surgical tool detection and $>$1pp (mAP) in surgical action recognition. Our proposed augmentation achieved comparable results in phase recognition, and improved $>$1pp (mAP) in tool detection and $>$1pp (F1-score) in action recognition.
Pretrained weights. We compare the effects of employing pretrained weights from ImageNet in the ConvNeXt-L backbone with those of using Surg-3M. As shown in Tables 3, 4, and 5, SurgFM (which also has a ConvNeXt-L backbone, but is pretrained on Surg-3M) outperformed the ConvNeXt-L backbone pretrained on ImageNet in all downstream tasks and datasets. This finding suggests that SurgFM serves as a superior starting point for surgery-specific downstream task fine-tuning compared to models pretrained on general-purpose datasets like ImageNet. This is clearly shown in the surgical tool presence detection experiments (Table 4) on the GraSP dataset, where SurgFM outperformed a ConvNeXt-L pretrained on ImageNet by more than 4pp in mAP.
6.2 Comparison with SotA
The performance of our SurgFM compared to the state of the art in each downstream task and dataset is shown in Fig. 1 and Tables 3, 4, and 5.
Table 5: Performance evaluation on surgical action recognition. Two datasets are used: CholecT50 [42] and SAR-RARP50 [48]. The metrics are computed as described in Sec. 5.1. The evaluation protocol and the meanings of G, S, and F are explained in Sec. 5.2. The superscript denotes that the SurgFM backbone is frozen during fine-tuning. The subscript ± indicates that both past and future frames are used [48].
Method Type CholecT50 SAR-RARP50
mAP (%) Accuracy (%) F1@10 (%)
ConvNeXt-L± [36] G $62.9\pm 2.3$ $71.9$ $70.4$
Tripnet [43] S 58.8$\pm$3.1 - -
Rendezvous [44] $60.4\pm 2.8$
Attention Tripnet [44] $61.1\pm 2.0$
SummerLab-AI± [48] - $\mathbf{81.5}$ $\mathbf{84.1}$
Uniandes± [48] - $78.6$ $82.3$
MS-TCN± [48, 16] - $69.0$ $70.7$
LoViT± [48, 35] - $59.8$ $43.0$
SurgFM (ours) F $\mathbf{64.2\pm 1.7}$ $70.2$ $68.4$
${}_{\pm}^{*}$ (ours) $63.7\pm 1.7$ $72.4$ $74.3$
SurgFM± (ours) $\mathbf{64.2\pm 1.7}$ $74.1$ $74.3$
Surgical phase recognition. As shown in Table 3, SurgFM outperformed all specialist [54, 27, 52, 26, 62, 12, 20, 29, 30, 35] and foundation [49, 24, 5] models in Cholec80 ($>$3pp Jacc.), AutoLaparo ($>$8pp Jacc.), and M2CAI16 ($>$4pp Jacc.). These results demonstrate that SurgFM is able to effectively extract features that are useful for different surgical procedures, capitalizing on the diversity of surgery types present in the Surg-3M dataset.
In GraSP, SurgFM demonstrated performance comparable (with an mAP difference of $<$1pp) to that of the methods presented in [17, 55, 4]. However, these competing methods [17, 55, 4] utilize both past and future frames (offline), whereas our SurgFM is restricted to observing only the current and preceding frames, reflecting its real-time operational focus, and its broader range of application [35]. Despite being constrained by its limited temporal perspective, SurgFM achieves comparable performance with these more information-rich approaches.
Surgical tool presence detection. As shown in Table 4, SurgFM outperformed all specialist [54, 27, 44] and foundation [49] models in Cholec80, and achieved a comparable performance to [44, 4] in GraSP.
Surgical action recognition. As shown in Table 5, SurgFM outperformed all specialist model [43, 44] in CholecT50 ($>$3pp mAP), and achieved a competitive performance to [48] in SAR-RARP50 ($<$7pp difference in accuracy).
Table 6: Leaderboard for the proposed tasks on Surg-3M. Multi-label (35 classes) video classification of procedure types and binary classification of surgery types.
Method Procedure type Surgery type
mAP (%) F1 (%) Accuracy (%) F1 (%)
SlowFast [17] 22.0 23.9 88.5 87.5
TimeSformer [6] 42.1 37.5 93.2 92.7
MViTv2 [34] 49.5 41.8 95.8 94.6
Video Swin Transformer [37] 51.4 47.9 98.8 98.7
SurgFM-Vid (ours) $\mathbf{57.8}$ $\mathbf{49.3}$ $\mathbf{98.9}$ $\mathbf{98.9}$
Frozen SurgFM model. As shown in Tables 3, 4, and 5, the frozen version of SurgFM, denoted SurgFM, is indeed a foundation model. It works on all datasets and tasks without retraining. In Cholec80, SurgFM outperforms specialist SotA architectures such as EndoNet, PhaseNet, and SV-RCNet, OHFM, TeCNO. In AutoLaparo, SurgFM outperforms all the six specialist models we compared it with. In M2CAI16, SurgFM outperforms two specialist SotA models, PhaseNet and SV-RCNet. In SAR-RARP50, SurgFM outperforms MS-TCN and LoViT.
Data scaling. We investigated the extent to which annotation effort can be alleviated by leveraging our pretrained SurgFM model. To do so, we fine-tuned it on downstream tasks with 10%, 50%, and 100% of the annotated data (Fig. 5). With 50% of the data, the performance decreased a maximum of 4.4pp across all datasets and tasks except GraSP, which has a total of four videos for training at 100%, and where performance dropped by 9.6pp. The closest result was achieved in phase recognition (Cholec80), with SurgFM achieving an accuracy of 92.3% on 50% of the data, only 0.1pp below prior SotA with 100% of the data [35]. Also at 50%, our SurgFM outperformed prior SotA [35] in AutoLaparo at 100%. These results suggest that the features learned by SurgFM from Surg-3M are highly informative. Comparing our results with other foundation models reported in the literature, SurgFM achieved $>$6pp (F1-score) than Ramesh et al. [49] in phase recognition (Cholec80), despite using only 10% of the data compared to their 12.5%. Furthermore, SurgFM also outperformed Ramesh et al. by $>$6pp (mAP) on tool presence detection (Cholec80) at 10%.
6.3 Leaderboard
To establish a baseline for the novel tasks presented in Sec. 3, we conducted an evaluation of SotA approaches on Surg-3M (Table 6), which serves as a benchmark for future research endeavors. By leveraging the robust representations extracted from SurgFM, our video classification model, SurgFM-Vid (Sec. 4), demonstrated significant improvements ($>$6 mAP) over all the leading video classification models tested on the multi-label procedure type classification task. On the binary surgery type classification task, our approach achieved comparable accuracy to the top performer, the Video Swin Transformer [37]. The incorporation of frame-level features extracted by SurgFM, in conjunction with the strategic application of prior knowledge through weighted feature embeddings informed by frame typicality, leads to substantial performance gains, underscoring the value of SurgFM as a critical component in surgical video analysis tasks. Despite these improvements, our results also highlight the challenges associated with surgical procedure classification, underscoring the need for continued innovation in this area.
7 Conclusion
Figure 5: Data scaling. Results of SurgFM when fine-tuned on the downstream tasks with different percentages of annotated training samples. The radius of the circles, computed following $2\log(\sigma+1)$, represents the standard deviation ($\sigma$) of each metric estimated over five cross-validation folds. Prior SotA is computed using 100% of the available data. Best viewed online.
Building on a large corpus of publicly available surgical videos, this work has two main contributions: we have created a large-scale and diverse surgical video dataset, Surg-3M, and a foundation model, SurgFM. Our SurgFM outperformed specialist architectures on four datasets and three downstream tasks, demonstrating its potential as a valuable tool for computer vision research in surgery. Furthermore, our experimental results show that SurgFM can maintain its performance with significantly reduced data annotation needs (50% or even 10% for some tasks), highlighting the efficiency of our approach. This work showcases the value of Surg-3M and SurgFM as a foundation for future research in surgical computer vision. To further improve these contributions, we plan to address key limitations of our current approach, including adapting SurgFM to pixel-wise tasks and enhancing the self-distillation approach to better adapt to surgical data. Ultimately, this work highlights the importance of publicly available surgical videos in developing more effective models for computer vision research in surgery, and we believe that Surg-3M and SurgFM will serve as catalysts for future advancements in this field.
[…]
Supplementary Material
You are a highly knowledgeable assistant specializing in surgical procedures and medical terminology.
Your expertise includes identifying and categorizing surgical interventions based on clinical
descriptions and procedural contexts.
Here is a list of 35 possible surgical procedure types: pancreatectomy, pancreaticoduodenectomy,
splenectomy, ampullectomy, hepatectomy, nephrectomy, low anterior resection, colectomy,
abdominoperineal resection, pulmonary lobectomy, hartmanns, prostatectomy, gastric bypass,
duodenal switch, gastrectomy, small bowel resection, hernia repair, ulcer repair, cholecystectomy,
appendectomy, ileocolic resection, cecectomy, myomectomy, hysterectomy, nissen fundoplication,
adrenalectomy, thymectomy, rectopexy, adhesiolysis, esophagectomy, cystectomy, jejunostomy, ileorectal
anastomosis, kidney transplant, vaginectomy.
Based on the description of the surgical video: <video title>, determine the most likely procedure type
from the list. Focus on matching the description to the procedure type that best aligns with the
terminology and context provided.
Figure S1: ChatGPT prompt employed to match video titles to procedure types. The title of the video to be matched is inserted where the $<$video title$>$ tag is located.
Appendix A Surgical Background
In contrast to traditional open surgery (not the target of our work), which entails extensive tissue disruption via large incisions and relies on direct visual inspection by the surgeon (i.e., no cameras are used), modern surgical practices predominantly employ minimally invasive techniques (the target of our work). These techniques involve the insertion of slender instruments, including one that is a camera, into the patient’s body through small incisions.
To control the instruments from outside the patient’s body, there are two variants: robotic-assisted surgery (called “robotic” in our manuscript) and non-robotic laparoscopy (called “non-robotic” in our manuscript). Robotic-assisted surgery involves a surgeon sitting on a console and controlling (with joysticks) several robotic arms that hold and steer the instruments inside the patient. In contrast, non-robotic (also known as traditional or conventional) laparoscopy requires the surgeon to directly hold and steer the laparoscopic hand-held instruments (i.e., no robotic arms involved).
Appendix B Details of Data Curation
This section provides details of the implementation of our data curation pipeline.
B.1 Video Classification
Video summarization. To improve the efficiency of annotating surgical and non-surgical videos, we obtained 4$\times$4-image video storyboards (i.e. a single image containing a collage of key video frames) for all the collected videos using the method described in [18]. Storyboards enabled us to quickly determine whether a video contained substantial surgical footage, thereby avoiding the need for complex analysis on the entire sequence.
Video storyboard classification. We manually annotated a dataset comprising 2160 surgical and 1910 non-surgical storyboards. The annotation criterion adopted for labeling a storyboard as surgical was that at least 50% of the key frames contained shots from a surgical camera; specifically, open surgery videos were categorized as non-surgical. To classify the rest of the collected videos, we trained a ResNet18 [22]. To ensure the accuracy of the inference results, we manually reviewed all the videos classified as surgical.
Performance of video storyboard classification models. We trained five video storyboard classification models (different data splits) to categorize the videos as either surgical or non-surgical. Each fold was split into training, validation, and testing sets with ratios of 0.8, 0.1, and 0.1, respectively. The average F1-score of the video storyboard classification models was $94.75\%\pm 1.1$. The results for each cross-validation fold are shown in Table S1.
Table S1: Performance of the video storyboard classification models across five folds.
Fold F1-score (%) Accuracy (%) Precision (%) Recall (%)
Fold 0 94.42 94.50 95.88 93.00
Fold 1 93.60 93.50 92.23 95.00
Fold 2 93.89 93.75 91.87 96.00
Fold 3 96.50 96.50 96.50 96.50
Fold 4 95.36 95.50 98.40 92.50
Average 94.75 94.75 94.98 94.60
Std Dev 1.10 1.16 2.66 1.57
B.2 Video Selection and Trimming
Frame classification. A ResNet18 [22] was trained for the surgical/non-surgical video frame classification task. To produce the annotations, the videos were sampled at one frame per second (fps). We annotated 7967 frames, 5481 of which turned to be surgical, and 2486, non-surgical.
Video trimming. Most online videos contain introductory and conclusion slides. Experimentally, we found that the start and end of the surgical footage can be reliably identified by finding the first and last three consecutive frames classified as surgical by our surgical frame classifier (sampling the video at 1 fps). Therefore, this is the approach we used to discard the non-surgical parts at the beginning and end of the collected videos. The resulting videos were manually quality controlled.
Performance of frame classification models. We trained five frame classification models to classify a video frame as either surgical or non-surgical. Each model of the five was trained in a different training-validation-testing split of the data, with a split rate of 0.8, 0.1, and 0.1. As shown in Table S2, the average F1-score of the frame classification models was $95.64\%\pm 0.94$.
Table S2: Performance of frame classification models across five folds.
Fold F1-score (%) Accuracy (%) Precision (%) Recall (%)
Fold 0 94.65 94.85 98.37 91.21
Fold 1 96.43 96.48 97.93 94.98
Fold 2 96.53 96.61 98.69 94.47
Fold 3 94.44 94.60 97.33 91.71
Fold 4 96.15 96.23 98.17 94.22
Average 95.64 95.75 98.10 93.32
Std Dev 0.94 0.88 0.49 1.39
B.3 Video Annotation
For the surgery type, a video is considered to be robotic if the video title includes any of the following keywords: Robotic, Robot, Robo, Hugo, Versius, Senhance, Telerobotic, Console, and da Vinci. The search for these terms in the video titles was case-insensitive. The remaining videos, with titles that do not include any of these keywords, are manually verified to ensure that they are manual surgical procedures. For the surgical procedure type, we cross-reference the video titles with the predefined list of procedures. When exact title matches are not found, we leverage the capabilities of the ChatGPTv4 API to perform a more nuanced analysis, incorporating a customized prompt as shown in Fig. S1. All annotations are manually quality controlled.
After curation, the final Surg-3M dataset resulted in $4194$ surgical videos, 94% of which are $1280\times 720p$, the remaining $6\%$ have varied resolutions, with a minimum of $640\times 480$ pixels.
Appendix C SurgFM Pretraining: Implementation Details
This section provides an overview of the pretraining process for the SurgFM model, and the image preprocessing pipeline employed.
SurgFM pretraining. For the pretraining of SurgFM, we utilized eight NVIDIA V100 GPUs (32 GB) with a batch size of 24 per GPU. We employed AdamW as the optimizer with a teacher temperature of 0.04, fp16 precision, an initial learning rate of 5e-4 (after warm-up), and a minimum learning rate of 1e-6. The model was trained for 60 epochs, including 10 warm-up epochs. This number of epochs enables the model to converge and stabilize its training loss on our dataset and pretext task. The model with the lowest training loss was selected as the final model.
Figure S2: Image preprocessing in SurgFM. We automatically detect and crop out the user interface information from the surgical frames.
Image preprocessing. We utilize our trained non-surgical bounding-box detection model to identify and obliterate UI elements, as demonstrated in Fig. S2.
Performance of non-surgical content detection models. We trained five non-surgical content detection models to detect and obliterate non-surgical regions in surgical video frames. Each model was trained on a different training-validation-testing split of the data, with a split ratio of 0.8, 0.1, and 0.1, respectively. The average mAP50 of the five models was $79.29\%\pm 3.4$, and the average mAP50-95 was $66.18\%\pm 2.98$. The results for all the cross-validation folds are shown in Table S3.
Table S3: Performance of bounding box detection models across five folds.
Fold Images Instances Box Precision (%) Box Recall (%) mAP50 (%) mAP50-95 (%)
Fold 0 272 447 74.79 77.63 75.50 62.00
Fold 1 272 438 70.73 77.40 76.36 61.79
Fold 2 272 509 83.34 79.60 84.89 68.56
Fold 3 272 422 75.27 83.69 80.31 68.98
Fold 4 272 496 81.32 81.45 82.58 68.65
Average 77.09 79.94 79.29 66.18
Std Dev 4.41 2.33 3.40 2.98
Appendix D Downstream Tasks: Data Splits and Fine-tuning
In this section, we provide the details of the datasets used to evaluate each downstream task along with their corresponding data splits.
Surgical phase recognition. For M2CAI16, we follow the data splits in [26, 54, 30], dividing the dataset into 27 training videos and 14 testing videos, respectively. For Cholec80, we adopt the data splits in [54, 12, 35, 27], allocating 40 videos for training, eight for validation, and 32 for testing. For AutoLaparo, we follow the standard split of ten training videos, four for validation, and seven for testing [57]. For GraSP, we adopt the data splits specified in [4], with four videos for training, four for validation, and five for testing.
Surgical tool presence detection. The data splits for Cholec80 and GraSP are identical to those employed in the surgical phase recognition task.
Surgical action recognition. For CholecT50, we follow the standard 5-fold data splits as proposed in [42]. For SAR-RARP50 [48] we use its recommended split of 27 videos for training, 13 for validation, and 10 for testing.
Fine-tuning on Downstream tasks. We used the teacher model from SurgFM as our backbone and employed three different protocols for downstream task fine-tuning. The images were resized to $448\times 448$ for all downstream tasks. The downstream models with linear heads were trained with a batch size of 112, an initial learning rate (LR) of 1e-4, the AdamW optimizer, and cross-entropy loss. An early stopping criterion was applied, stopping training after 10 epochs if the validation loss shows no improvement.
For the surgical tool presence detection tasks and the CholecT50 action recognition task, we used a linear head. For the surgical phase recognition tasks, we utilized a TCN [12] head and followed the two-stage training approach outlined in TeCNO [12]. In the first stage, we trained the backbone with a linear head to perform frame-wise classification without temporal context. Then, we employed the TCN head to incorporate temporal information of the extracted features for predictions. The default TCN configuration was used, consisting of two stages, each containing nine layers. The TCN head was trained using the Adam optimizer with an initial learning rate of 5e-3 for 70 epochs to ensure sufficient training. The checkpoint with the lowest validation loss was selected as the optimal checkpoint. For the SAR-RARP50 [48] action recognition task, we trained our model in two modes: online, where the SurgFM teacher model was used as backbone with a TCN head that could see only the information from previous frames; and offline, where the SurgFM teacher model was paired with an MS-TCN head [33] to incorporate both previous and future frame information for a fair comparison with the methods in [48]. For the TCN head, we maintained the same model configuration and hyperparameters as in the surgical phase recognition tasks. For the MS-TCN head, we adopted a 3-stage, 11-layer configuration, following the basic model setup in [33]. The model was trained using the Adam optimizer with an initial learning rate of 5e-4 for 70 epochs.
Data scaling. In Sec. 6 of the manuscript, we present the results of the data scaling experiments. For each scale (10%, 25%, 50%), we conducted a 5-fold cross-validation, with each fold randomly selected from the training dataset. The model configurations and hyperparameters were kept consistent across different data scales using the same settings as those used for the full training dataset.
Appendix E Evaluation of SurgFM-Vid
To evaluate the performance of our video classifier, SurgFM-Vid, we have split the Surg-3M dataset by videos, with 3369 videos for training and 825 videos for testing. Our vanilla SurgFM was trained on all the Surg-3M frames, therefore, to evaluate the video classification performance fairly, we have trained a new SurgFM model from scratch using only the frames of the 3369 videos in the training set that are accessible to all other classifiers in the comparison (i.e., 2.6 million frames instead of the usual 3.3 million frames in Surg-3M).
Appendix F Dataset and Licensing
We adhere to the same practices as other datasets created from YouTube and various sources on the Web, such as ImageNet [13], Kinetics [9], YouTube-VIS [61], YouTube-8M [1], Insect-1M [41], Moments-in-Time [40], Tai-Chi-HD [51], HD-Villa-100M [60], and AVSpeech [15]. Specifically, we provide the code to download and generate the Surg-3M dataset, a list of links to the original YouTube videos, and the corresponding annotations. Researchers working in academic institutions can request direct access to the full Surg-3M dataset in LMDB format for non-commercial purposes. We will also provide an online form for the YouTube video authors to opt out of our Surg-3M dataset. The Surg-3M dataset is provided under the Creative Commons Attribution 4.0 International (CC BY 4.0) license.

### v2

@@
 To address these constraints, a new dataset called Surg-3M has been compiled using a novel aggregation pipeline that collects high-resolution videos from online sources.
-Featuring an extensive collection of over 4K surgical videos and more than 3 million high-quality images from multiple procedure types, Surg-3M offers a comprehensive resource surpassing existing alternatives in size and scope, including two novel tasks.
+Featuring an extensive collection of over 4K surgical videos totaling 938 hours of high-quality footage across multiple procedure types, Surg-3M offers a comprehensive resource surpassing existing alternatives in size and scope, including two novel tasks.
 To demonstrate the effectiveness of this dataset, we present SurgFM, a self-supervised foundation model pretrained on Surg-3M that achieves impressive results in downstream tasks such as surgical phase recognition, action recognition, and tool presence detection.
@@
 Building upon this data curation pipeline, we present Surg-3M, a dataset for computer vision applications in surgery.
-Our Surg-3M dataset is large and diverse, containing 4K surgical videos and 3M images (Table 1).
+Our Surg-3M dataset is large and diverse, comprising more than 4K surgical videos totaling 938 hours of high-quality footage (Table 1).
 To the best of our knowledge, Surg-3M is the largest open-access surgical video dataset to date, 18×18$\times$18 × larger than the largest published surgical dataset [54].
@@
 Table 1: Comparison of Surg-3M’s size to other existing public surgical datasets. The number of frames is obtained by sampling videos at one frame per second.
-Dataset Year # Procedure types # Videos # Frames
-M2CAI16 [54] $2016$ 1 41 $95$K
-Cholec80 [54] $2016$ 1 80 $183$K
-CholecT50 [44] $2022$ 1 50 $101$K
-AutoLaparo [57] $2022$ 1 21 $83$K
-SAR-RARP50 [48] $2023$ 1 50 $163$K
-GraSP [4] $2024$ $1$ $13$ $117$K
-Surg-3M (ours) $\mathbf{2025}$ $\mathbf{35}$ $\mathbf{4194}$ $\mathbf{3}$M
+Dataset # Procedures # Videos # Hours # Frames # Frames (1 fps)
+M2CAI16 [54] 1 41 $26$ $2.4$M $95$K
+Cholec80 [54] 1 80 $51$ $4.6$M $183$K
+CholecT50 [44] 1 50 $28$ $2.5$M $101$K
+AutoLaparo [57] 1 21 $23$ $2.1$M $83$K
+SAR-RARP50 [48] 1 50 $45$ $9.8$M $163$K
+GraSP [4] $1$ $13$ $33$ $3.5$M $117$K
+Surg-3M (ours) $\mathbf{35}$ $\mathbf{4194}$ $\mathbf{938}$ $\mathbf{85}$M $\mathbf{3.4}$M
 2 Related Work
@@
 To develop a robust surgical foundation model, it is essential to create a comprehensive, large-scale dataset of diverse surgical images that covers the whole spectrum of minimally invasive surgical procedures executed within the operating room.
-To fulfill this need we propose Surg-3M, a dataset that contains 4K videos and 3M images corresponding to 35 different surgical procedures, including videos from robotic-assisted surgeries and non-robotic traditional laparoscopies (more details on these surgical techniques are provided in Sec.
+To fulfill this need we propose Surg-3M, a dataset comprising more than 4K videos totaling 938 hours of surgical footage covering 35 distinct procedures, including videos from robotic-assisted surgeries and non-robotic traditional laparoscopies (more details on these surgical techniques are provided in Sec.
 A of the supplementary material).
 Ground truth and proposed new tasks.
-Each of the 4K surgical videos is annotated for two tasks: 1) multi-label procedure type classification, where the objective is to identify the procedure[s] performed in an intervention’s video from among 35 possible options (Fig.
+Each of the 4194 surgical videos is annotated for two tasks: 1) multi-label procedure type classification, where the objective is to identify the procedure[s] performed in an intervention’s video from among 35 possible options (Fig.
 2); and 2) binary surgery type classification, where the goal is to indicate whether a surgery is conducted robotically or non-robotically.
@@
 Scenes in surgical videos unfold at a relatively slow pace, as surgeons must ensure patient safety.
-Consequently, our 3M image dataset was compiled by sampling our dataset of 4194 curated videos at 1 fps, which is standard practice in the field of surgical computer vision [54, 57, 44, 48, 53].
+Consequently, we extract frames from Surg-3M at 1 fps for our foundation model pretraining, following the standard practice in the field of surgical computer vision [54, 57, 44, 48, 53].
 In the following sections we detail how we collected, curated, and annotated these videos.

### v3

@@
-Surg-3M: A Dataset and Foundation Model for Perception in Surgical Settings
+LEMON: A Large Endoscopic MONocular Dataset and Foundation Model for Perception in Surgical Settings
 Abstract
-Advancements in computer-assisted surgical procedures heavily rely on accurate visual data interpretation from camera systems used during surgeries.
-Traditional open-access datasets focusing on surgical procedures are often limited by their small size, typically consisting of fewer than 100 videos with less than 100K images.
-To address these constraints, a new dataset called Surg-3M has been compiled using a novel aggregation pipeline that collects high-resolution videos from online sources.
-Featuring an extensive collection of over 4K surgical videos totaling 938 hours of high-quality footage across multiple procedure types, Surg-3M offers a comprehensive resource surpassing existing alternatives in size and scope, including two novel tasks.
-To demonstrate the effectiveness of this dataset, we present SurgFM, a self-supervised foundation model pretrained on Surg-3M that achieves impressive results in downstream tasks such as surgical phase recognition, action recognition, and tool presence detection.
-Combining key components from ConvNeXt, DINO, and an innovative augmented distillation method, SurgFM exhibits exceptional performance compared to specialist architectures across various benchmarks.
-Our experimental results show that SurgFM outperforms state-of-the-art models in multiple downstream tasks, including significant gains in surgical phase recognition (+8.9pp, +4.7pp, and +3.9pp of Jaccard in AutoLaparo, M2CAI16, and Cholec80), action recognition (+3.1pp of mAP in CholecT50) and tool presence detection (+4.6pp of mAP in Cholec80).
-Moreover, even when using only half of the data, SurgFM outperforms state-of-the-art models in AutoLaparo and achieves state-of-the-art performance in Cholec80.
-Both Surg-3M and SurgFM have significant potential to accelerate progress towards developing autonomous robotic surgery systems.
-Our dataset, models, and code are publicly available at https://github.com/visurg-ai/surg-3m.
+Traditional open-access datasets focusing on surgical procedures are often limited by their small size, typically consisting of fewer than 100 videos and less than 30 hours of footage, which leads to poor model generalization.
+To address this constraint, a new dataset called LEMON has been compiled using a novel aggregation pipeline that collects high-resolution videos from online sources.
+Featuring an extensive collection of over 4K surgical videos totaling 938 hours (85 million frames) of high-quality footage across multiple procedure types, LEMON offers a comprehensive resource surpassing existing alternatives in size and scope, including two novel downstream tasks.
+To demonstrate the effectiveness of this diverse dataset, we introduce LemonFM, a foundation model pretrained on LEMON using a novel self-supervised augmented knowledge distillation approach.
+LemonFM consistently outperforms existing surgical foundation models across four downstream tasks and six datasets, achieving significant gains in surgical phase recognition (+9.5pp, +9.4pp, and +8.4pp of Jaccard in AutoLaparo, M2CAI16, and Cholec80), surgical action recognition (+4.4pp of mAP in CholecT50), surgical tool presence detection (+5.3pp and +10.2pp of mAP in Cholec80 and GraSP), and surgical semantic segmentation (+8.3pp of mDice in CholecSeg8k).
+LEMON and LemonFM will serve as foundational resources for the research community and industry, accelerating progress in developing autonomous robotic surgery systems and ultimately contributing to safer and more accessible surgical care worldwide.
+Our dataset, models, and code are publicly available at https://github.com/visurg-ai/LEMON.
 1 Introduction
-Figure 1: Performance comparison between our foundation model, SurgFM, and the state-of-the-art (SotA) models.
-Our evaluation focuses on three surgical downstream tasks and six datasets.
-SurgFM results are shown in bold.
-Axis labels are written in regular font.
-For a detailed breakdown of the results, see Tables 3, 4, and 5.
-Figure 2: Diversity and procedure prevalence in Surg-3M.
-Representative samples from various procedures, demonstrating the diverse range of cases in our curated dataset (left, right).
-Distribution of surgical frames by procedure type (center).
-Autonomous surgery has long been a promising area of research, with the potential to revolutionize the way surgical procedures are performed [25].
-To achieve this goal, significant advancements in perception and computer vision are necessary, allowing robotic systems to accurately interpret and respond to the complexities of surgical environments [38].
-However, training models for surgical perception is far from trivial due to the limited availability and high quality requirements of annotated data [59, 39, 45].
-The complexities of medical data and stringent regulations governing patient privacy create significant hurdles for collecting and annotating relevant information [19, 2].
-The cumulative effect of all these challenges has resulted in significantly limited computer vision datasets for surgical applications, with many containing fewer than 10K images [50].
-These datasets, which may suffice for validation on specific computer vision tasks and surgical procedures, fall short in providing an adequate foundation for training models that can generalize effectively across diverse surgical settings and datasets.
-Recent advances in self-supervised learning have opened up new research avenues, pretrained foundation models can significantly expedite progress, potentially reducing the need for annotated data by utilizing large-scale unlabeled datasets [49].
-Despite this progress, ongoing data collection challenges persist, motivating the exploration of two key aspects: the sufficiency of publicly available surgical videos as a foundation for building a comprehensive dataset, and the potential of self-supervised models trained on this dataset to achieve state-of-the-art performance in surgical downstream tasks.
-By addressing these aspects, we aim to provide valuable insights into the feasibility and efficacy of using public domain data for training machine learning models in the field of surgical video analysis.
+Figure 1: Performance comparison between LemonFM, our foundation model pretrained on LEMON, and other surgical foundation models.
+The numeric labels mark the axis ticks.
+The full results are available in Table 3 and 4.
 Figure 2: Data curation pipeline.
@@
 We further filter videos by requiring at least 90% of their frames to be surgical frames.
-(d) In the video preprocessing stage, we obliterate non-surgical frames.
+(d) In the video preprocessing stage, we obliterate non-surgical frames and non-surgical regions within surgical frames.
 (e) During the video annotation phase, we utilize the video title as a primary cue to determine the procedure and surgery type.
-In cases where the video titles do not explicitly include any of our procedure name keywords (Fig.
-2), we employ ChatGPT to match the titles with surgical procedure types.
-The set of final videos, and their corresponding labels, are manually quality controlled.
+In cases where the video titles do not explicitly include any of our procedure name keywords, we employ ChatGPT to match the titles with surgical procedure types.
+Both the final videos and their corresponding labels are manually quality controlled.
+Autonomous surgery has long been a promising area of research, with the potential to revolutionize the way surgical procedures are performed [24].
+It promises safer and more consistent operations, fewer complications, shorter operative times, lighter workloads for surgeons, and wider access to high‑quality care.
+To achieve this goal, significant advancements in perception and computer vision are necessary, allowing robotic systems to accurately interpret and respond to the complexities of surgical environments [38].
+However, training models for surgical perception is far from trivial due to the limited availability and high quality requirements of annotated data [54, 40, 44].
+The complexities of medical data and stringent regulations governing patient privacy create significant hurdles for collecting and annotating relevant information [17, 2].
+The cumulative effect of all these challenges has resulted in significantly limited computer vision datasets for surgical applications, with many containing fewer than 10K images [47].
+These datasets, which may suffice for validation on specific computer vision tasks and surgical procedures, fall short in providing an adequate foundation for training models that can generalize effectively across diverse surgical settings and datasets.
+Recent advances in self-supervised learning have opened up new research avenues.
+Pretrained foundation models can significantly expedite progress, potentially reducing the need for annotated data by utilizing large-scale unlabeled datasets [46].
+Despite this progress, ongoing data collection challenges persist, motivating the exploration of two key aspects: the sufficiency of publicly available surgical videos as a foundation for building a comprehensive dataset, and the potential of self-supervised models trained on this dataset to achieve state-of-the-art performance in surgical downstream tasks.
+By addressing these aspects, we aim to provide valuable insights into the feasibility and efficacy of using public domain data for training machine learning models in the field of surgical vision analysis.
+Table 1: Comparison of LEMON’s size to other existing public surgical datasets.
+Dataset Procedure Video Hour Frame Frame (1 fps)
+M2CAI16 1 41 262626 2.42.42.4M  [^95] K
+Cholec80 1 80 515151 4.64.64.6M 183183183K
+CholecT50 1 50 282828 2.52.52.5M 101101101K
+AutoLaparo 1 21 232323 2.12.12.1M  [^83] K
+GraSP 111 131313 333333 3.53.53.5M 117117117K
+GenSurgery 282828 318231823182 680680680  [^70] M 2.42.42.4M
+SurgeNetXL 232323 325332533253 576576576  [^52] M 2.12.12.1M
+LEMON $\mathbf{35}$ $\mathbf{4194}$ $\mathbf{938}$ $\mathbf{85}$M $\mathbf{3.4}$M
+Figure 3: Comparison between LEMON and other surgical datasets sourced from the web.
+Video segments from each dataset illustrating the data curation differences (left).
+Procedures which are not covered by both GenSurgery and SurgeNetXL or any other public dataset (right).
 Contributions.
-We propose a novel data curation pipeline (Fig.
-3) for collecting videos of selected surgical procedures, capturing exclusively surgical footage while discarding non-surgical content.
-Building upon this data curation pipeline, we present Surg-3M, a dataset for computer vision applications in surgery.
-Our Surg-3M dataset is large and diverse, comprising more than 4K surgical videos totaling 938 hours of high-quality footage (Table 1).
-To the best of our knowledge, Surg-3M is the largest open-access surgical video dataset to date, 18×18$\times$18 × larger than the largest published surgical dataset [54].
-To model the features of surgical video frames, we provide SurgFM, a pretrained model based on the ConvNeXt architecture [36] and DINO [8].
-Additionally, we propose a novel augmented distillation method (Fig.
-4) designed to enhance the local-to-global self-distillation mechanism described in  [8].
-SurgFM outperformed the current state-of-the-art models (Fig.
-1) on Cholec80 [54], M2CAI16 [54, 53], AutoLaparo [57], and CholecT50 [44].
-Additionally, it helps to significantly reduce the annotation effort (Fig.
-5).
-Building on our Surg-3M annotations, we introduce two new supervised learning applications: multi-label surgical procedure classification and surgery type classification.
-Leveraging our image foundation model, SurgFM, we propose SurgFM-Vid, a video classifier that achieves top performance on these new tasks, serving as a baseline for future research.
-Table 1: Comparison of Surg-3M’s size to other existing public surgical datasets. The number of frames is obtained by sampling videos at one frame per second.
-Dataset # Procedures # Videos # Hours # Frames # Frames (1 fps)
-M2CAI16 [54] 1 41 $26$ $2.4$M $95$K
-Cholec80 [54] 1 80 $51$ $4.6$M $183$K
-CholecT50 [44] 1 50 $28$ $2.5$M $101$K
-AutoLaparo [57] 1 21 $23$ $2.1$M $83$K
-SAR-RARP50 [48] 1 50 $45$ $9.8$M $163$K
-GraSP [4] $1$ $13$ $33$ $3.5$M $117$K
-Surg-3M (ours) $\mathbf{35}$ $\mathbf{4194}$ $\mathbf{938}$ $\mathbf{85}$M $\mathbf{3.4}$M
+To address the scarcity of high‑quality surgical data and the practical difficulty of collecting it, we propose a novel multi-stage data curation pipeline designed specifically for online surgical videos (Fig.
+2).
+In contrast to previous efforts that relied on limited, manually curated validation processes with datasets containing non-surgical content, our pipeline employs a systematic approach to filter out non-surgical videos, remove non-surgical content from surgical videos, and verify the curated surgical videos and their annotations.
+Building upon this data curation pipeline, we present LEMON, a large-scale dataset comprising more than 4K surgical videos totaling 938 hours (85 million frames) of high-quality footage (Table 1).
+Additionally, we propose two novel supervised video classification tasks based on the accompanying annotations: multi-class surgical procedure classification and binary surgery type classification, and establish a leaderboard for them.
+To the best of our knowledge, LEMON is the largest open access surgical dataset to date, both in total frames and videos and in the breadth of distinct surgical procedures it includes.
+To model the features of surgical frames, we introduce LemonFM, a foundation model pretrained on LEMON that leverages self-supervised learning with our proposed augmented knowledge distillation method.
+We evaluated LemonFM on six datasets: AutoLaparo [52], Cholec80 [51], M2CAI16 [51], GraSP [4], CholecT50 [43], and CholecSeg8k [23], covering four tasks including phase recognition, tool presence detection, action recognition, and semantic segmentation.
+Under both linear probing and full fine-tuning protocols, LemonFM consistently surpasses existing surgical foundation models and task-specific architectures (Fig.
+1, Table 2, 3, 4).
+Moreover, LemonFM retains these advantages even when fine‑tuned with only half of the labeled data.
+These results showcase LemonFM’s strong generalization and data efficiency over diverse surgical scenarios and underscore the value of LEMON as a broadly useful resource for the community.
+We provide open access to our dataset, models, and code for data curation and training.
 2 Related Work
 Surgical datasets.
-Previous works have released open-access datasets aimed at addressing specific surgical downstream tasks and procedures (Table 1).
-Despite their valuable contributions, these open-access datasets suffer from a narrow scope of surgical procedures and limited dataset sizes.
-Besides these open-access datasets, there are several private surgical datasets, including SVL [63], with 1.4K videos, and SAGES CVS [47], with 1K videos.
-However, the restricted accessibility of these datasets constrains their utility and hinders their potential for more widespread use.
-To overcome these limitations, our work introduces Surg-3M.
-Compared with Cholec80 [54], the largest open-access surgical dataset, Surg-3M provides 18$\times$ more surgical images and 35$\times$ the surgical procedure types.
+Previous works have released open-access datasets aimed at addressing specific surgical vision tasks and procedures (Table 1).
+Despite valuable, these datasets suffer from a narrow scope of surgical procedures and limited dataset sizes.
+GenSurgery [48] and SurgeNetXL [25] expand procedural coverage by compiling public videos, but the lack of a curation step to remove the non-surgical content from the public videos (Fig.
+3) might introduce spurious features that can confound discriminative models.
+Additionally, several private datasets exist, such as SVL [57] and SAGES CVS [45], but their restricted access limits their utility and broader adoption.
+To overcome these limitations, we introduce a novel data curation pipeline that yields LEMON, the largest open-access surgical dataset to date, surpassing previous datasets in both the number of frames and videos and the diversity of distinct surgical procedures represented.
 Self-supervised learning.
@@
 Recently, Ramesh et al.
-[49] investigated the effectiveness of different SSL approaches for surgical applications, including MoCov2 [11], SwAV [7], SimCLR [10], and DINO [8].
-Their experiments showed that DINO consistently outperformed other SSL methods on both surgical phase recognition and surgical tool presence detection tasks, highlighting its potential in surgical applications.
+[46] investigated the effectiveness of different SSL approaches for surgical applications.
+Their experimental results show that DINO [8] outperformed MoCov2 [11], SwAV [7], and SimCLR [10] in surgical applications.
 As part of a broader effort to advance self-supervised surgical models, Hirsch et al.
-[24] built a private surgical dataset and trained Masked Siamese Networks (MSN) [3] on this dataset, achieving near-state-of-the-art results.
-While the research presented in this study shows promising results, its broader applicability and reproducibility are somewhat constrained by the lack of public availability of the underlying dataset, as well as a relatively narrow evaluation scope that only considers performance on two specific task-dataset pairs: surgical phase recognition on Cholec80 [54], and polyp characterization on PolypSet [32].
+[22] built a private surgical dataset and trained a Masked Siamese Networks (MSN) [3] on it.
+Despite encouraging results, the work’s wider applicability and reproducibility are limited by its reliance on private data and a relatively narrow evaluation scope: surgical phase recognition on Cholec80 and polyp characterization on PolypSet [33].
 Similarly, in Endo-FM [53], the authors proposed to train a student-teacher foundation model on a composite dataset that combined private and public data sources.
-Although showing encouraging results, the approach is also of limited accessibility.
+Although showing promising results, the approach is also of limited accessibility.
 More recently, Batic et al.
-[5] created a dataset of 70K images merging nine public datasets.
+[5] created a dataset by merging nine public datasets.
 They trained a model based on masked autoencoders (MAE) [21], but showed limited performance, likely due to their reduced dataset size.
-[…]
-Method Phase recognition (Cholec80) Tool detection (Cholec80) Action recognition (CholecT50)
-Accuracy (%) F1-score (%) mAP (%) F1-score (%) mAP (%) F1-score (%)
-DINO (frozen) [8] $69.8\pm 1.5$ $62.5\pm 1.5$ $67.8\pm 3.1$ $59.3\pm 4.0$ $36.1\pm 1.4$ $28.9\pm 2.7$
-DINOv2 (frozen) [46] $67.0\pm 1.7$ $60.3\pm 1.8$ $56.1\pm 3.4$ $41.2\pm 2.6$ $27.2\pm 2.2$ $17.2\pm 0.8$
-MAE (frozen) [23] $55.2\pm 1.4$ $49.6\pm 1.2$ $40.4\pm 1.4$ $22.9\pm 0.9$ 20.7 $\pm$ 1.3 13.4 ±$\pm 1.3$
-[…]
-[…]
-[…]
-3 Proposed Surg-3M Dataset
+3 Proposed LEMON Dataset
 To develop a robust surgical foundation model, it is essential to create a comprehensive, large-scale dataset of diverse surgical images that covers the whole spectrum of minimally invasive surgical procedures executed within the operating room.
-To fulfill this need we propose Surg-3M, a dataset comprising more than 4K videos totaling 938 hours of surgical footage covering 35 distinct procedures, including videos from robotic-assisted surgeries and non-robotic traditional laparoscopies (more details on these surgical techniques are provided in Sec.
-A of the supplementary material).
-Ground truth and proposed new tasks.
-Each of the 4194 surgical videos is annotated for two tasks: 1) multi-label procedure type classification, where the objective is to identify the procedure[s] performed in an intervention’s video from among 35 possible options (Fig.
-2); and 2) binary surgery type classification, where the goal is to indicate whether a surgery is conducted robotically or non-robotically.
-The rationale for these annotations is two-fold.
-Firstly, it allows us to validate the comprehensiveness of our dataset by ensuring that it encompasses a broad spectrum of surgical procedures.
-Secondly, it facilitates the development of new methods that use this information as a key input (e.g., our SurgFM foundation model, video classification/retrieval, video similarity measurement, procedure-specific estimation of surgical remaining time).
-Scenes in surgical videos unfold at a relatively slow pace, as surgeons must ensure patient safety.
-Consequently, we extract frames from Surg-3M at 1 fps for our foundation model pretraining, following the standard practice in the field of surgical computer vision [54, 57, 44, 48, 53].
-In the following sections we detail how we collected, curated, and annotated these videos.
-An overview of the curation pipeline is shown in Fig.
-3.
+To fulfill this need, we propose LEMON, a dataset comprising more than 4K videos totaling 938 hours of surgical footage that covers 35 distinct procedures, including videos from robotic-assisted surgeries and non-robotic traditional laparoscopies (see more details on these surgical techniques in the supplementary material).
 3.1 Dataset Curation
+We detail how we collected, curated, and annotated these videos. An overview of the pipeline is shown in Fig. 2.
 Video collection.
+We first compiled a list of 35 distinct minimally invasive surgical procedure types [39] (see Fig.
+1 in supplementary material).
 Inspired by works in other fields [14, 9, 56, 1, 41], we collected our videos from YouTube.
-For each of the $35$ procedures in Fig.
-2, we searched for videos of their robotic and non-robotic variants using the strings “robotic $<$procedure type$>$” and “laparoscopic $<$procedure type$>$”, respectively.
-We run the video collection process until we had $500$ different videos collected for each procedure leading to a total of $\approx$18K raw videos.
+For each of the 353535 procedures, we searched for videos of their robotic and non-robotic variants using the strings “robotic $<$procedure type$>$” and “laparoscopic $<$procedure type$>$”, respectively.
+We run the video collection process until we had 500500500 different videos collected for each procedure leading to a total of $\approx$18K raw videos.
 Video classification.
@@
 To avoid including non-surgical videos in our dataset, we classified the videos as surgical and non-surgical (Fig.
-3b).
-This pipeline consisted of three stages.
-First, a dataset of 2160 surgical and 1910 non-surgical video storyboards (Fig.
-3b) was annotated.
-Second, a binary video storyboard classifier was trained on this dataset and run over all the raw videos.
+2b).
+First, we created 4$\times$4-image video storyboards (i.e., a single image containing 16 key video frames) for each video using the method described in [16].
+Second, a dataset of 2160 surgical and 1910 non-surgical video storyboards (Fig.
+2b) was annotated.
+Third, a binary video storyboard classifier (i.e., a ResNet18 [20]) was trained on this dataset and run over all the raw videos.
 Finally, manual verification was performed on the videos classified as surgical.
 Video selection and trimming.
-A video classified as surgical may still contain non-surgical frames (e.g., presentation slides).
-Therefore, we trained a surgical frame classifier and used it to 1) trim the non-surgical content present at the beginning and end of videos, and 2) discard videos containing more than 10% of the frames estimated as non-surgical after trimming (Fig.
-3c).
+A video classified as surgical may still contain non-surgical frames (e.g., introductory and conclusion slides).
+Experimentally, we found that the start and end of the surgical footage can be reliably identified by finding the first and last three consecutive surgical frames (sampling the video at 1 fps).
+Thus, we manually annotated a dataset of 7967 frames (5481 surgical and 2486 non-surgical), trained a surgical frame classifier (i.e., a ResNet18) on it, and used this model to: 1) trim the non-surgical content present at the beginning and end of videos, and 2) discard videos containing more than 10% of the frames estimated as non-surgical after trimming (Fig.
+2c).
 Video preprocessing.
-We use our previously trained frame classifier to detect and obliterate those frames that are estimated to be non-surgical within the trimmed videos.
-After this process, we manually quality control the videos to verify that their frames are surgical.
+As the videos were from YouTube, we cannot assume that all frames within the trimmed videos are entirely free from non-surgical content (e.g., intraoperative out-of-body views, medical annotations, user-interface details, equipment logos).
+Therefore, we used our trained frame classifier to detect and obliterate the intraoperative non-surgical frames (Fig.
+2d).
+Additionally, we manually annotated 2719 surgical frames with 4584 non-surgical bounding-box instances and trained a YOLOv8 model [31] to detect and obliterate the non-surgical content in surgical frames.
+After this process, we manually quality controlled all the remaining curated videos.
 Video annotation.
-We cross-reference the video titles with the predefined list of procedures.
-When exact title matches are not found, we leverage the capabilities of the ChatGPTv4 API to perform a more nuanced analysis, incorporating a customized prompt as specified in Sec.
-B of the supplementary material.
-All annotations are manually reviewed and validated to ensure the accuracy of the labels assigned.
-4 Proposed SurgFM Foundation Model
-In this section, we describe how we build SurgFM, our foundation model trained on Surg-3M, and motivate its design choices.
-This section includes two ablation studies to identify the optimal self-distillation configuration and the most suitable backbone architecture for surgical videos; a description of our image preprocessing pipeline, and our proposed augmented distillation mechanism with its corresponding ablation study.
-The results of the ablation experiments are presented in Table 2.
-Self-supervised methods.
-We examine the impact of the different self-supervision mechanisms (Table 2).
-We evaluate DINO [8], DINOv2 [46], and MAE [23].
-For these experiments, we trained the models in their standard configurations, a ViT-B backbone [14] pretrained on ImageNet-1k [13] with frozen weights and a linear classification head.
-Stemming from these results, which align with those in the literature [49], DINO consistently outperforms DINOv2 and MAE on all downstream tasks, with surgical tool detection and action recognition showing a large margin.
-Therefore, we adopt DINO’s self-distillation [8] as our pretext task for self-supervision.
-Backbone models.
-As noted in [56], ViTs [14] face challenges to capture fine details, which is essential to solving tasks such as tool presence detection and surgical action recognition.
-Given this, we opted to evaluate the performance of ViT-L [14] against a modern convolutional architecture, ConvNeXt-L [36].
-Both were pretrained on Surg-3M and fine-tuned on Cholec80 [54] and CholecT50 [42], for surgical phase recognition, tool presence detection, and surgical action recognition.
-The results are shown in Table 2.
-ConvNeXt-L outperformed ViT-L, suggesting that convolutional networks are indeed better equipped to capture those fine image details that are relevant to solve surgical tasks.
-Therefore, we use ConvNeXt-L as the backbone of our SurgFM model.
-Image preprocessing pipeline.
-Surgical videos often exhibit superimposed manufacturer-specific user interface (UI) information generated by the software of the surgical stack.
-While certain researchers might derive value from analyzing these UI elements (e.g., some interfaces provide contextual information regarding the surgical instruments utilized in the video), a foundational model for surgery should prioritize learning correlations between tissue appearance and tool-tissue interactions.
-This focus is essential to avoid acquiring spurious correlations with manufacturer-specific UI elements.
-Therefore, we manually annotated 2719 surgical frames with 4584 non-surgical bounding-box instances and trained a YOLOv8 Nano model [31] to detect and crop out UI content (see Fig.
-S2 in supplementary material).
-[…]
-Method Type Cholec80 AutoLaparo M2CAI16 GraSP
-Accuracy (%) Precision (%) Recall (%) Jaccard (%) F1-score (%) Accuracy (%) Precision (%) Recall (%) Jaccard (%) Accuracy (%) Precision (%) Recall (%) Jaccard (%) mAP (%)
-ConvNeXt-L [36] G $91.4\pm 6.7$ $92.3\pm 3.8$ $90.5\pm 7.7$ $82.1\pm 7.1$ $88.6\pm 6.8$ $82.3\pm 9.7$ $78.3\pm 21.9$ $69.9\pm 31.9$ $60.5\pm 27.2$ $88.3\pm 10.0$ $89.4\pm 4.7$ $89.5\pm 3.3$ $78.4\pm 6.2$ $73.4$
-EndoNet [54] S $81.7\pm 4.2$ $73.7\pm 16.1$ $79.6\pm 7.9$ $76.5$
-MTRCNet-CL [27] $89.2\pm 7.6$ $86.9\pm 4.3$ $88.0\pm 6.9$ $87.4$
-PhaseNet [52] $78.8\pm 4.7$ $71.3\pm 15.6$ $76.6\pm 16.6$ $73.8$ $79.5\pm 12.1$ $64.1\pm 10.3$
-SV-RCNet [26] $85.3\pm 7.3$ $80.7\pm 7.0$ $83.5\pm 7.5$ $82.1$ 75.6 $64.0$ $59.7$ $47.2$ 81.7 $\pm$ 8.1 81.0 $\pm$ 8.3 81.6 $\pm$ 7.2 65.4 $\pm$ 8.9 -
-OHFM [62] $87.3\pm 5.7$ $67.0\pm 13.3$ $85.2\pm 7.5$ $68.8\pm 10.5$
-TeCNO [12] $88.6\pm 7.8$ $86.5\pm 7.0$ $87.6\pm 6.7$ $75.1\pm 6.9$ $77.3$ $66.9$ $64.6$ $50.7$ $86.1\pm 10.0$ $85.7\pm 7.7$ $88.9\pm 4.5$ $74.4\pm 7.2$
-TMRNet [29] $90.1\pm 7.6$ $90.3\pm 3.3$ $89.5\pm 5.0$ $79.1\pm 5.7$ $78.2$ $66.0$ $61.5$ $49.6$
-AVT [20] - - - - - $77.8\pm 9.4$ $68.0$ $62.2$ $50.7$
-Trans-SVNet [30] $90.3\pm 7.1$ $90.7\pm 5.0$ $88.8\pm 7.4$ $79.3\pm 6.6$ $78.3$ $64.2$ $62.1$ $50.7$ $87.2\pm 9.3$ $88.0\pm 6.7$ $87.5\pm 5.5$ $74.7\pm 7.7$
-LoViT [35] $92.4\pm 6.3$ $89.9\pm 6.1$ $90.6\pm 4.4$ $81.2\pm 9.1$ $81.4\pm 7.6$ $85.1$ $65.9$ $55.9$
-SlowFast± [17] - - - - - - - - - - - - - $70.1$
-TAPIR± [55] - - - - - - - - - - - - - $74.6$
-TAPIS± [4] - - - - - - - - - - - - - $\mathbf{76.1}$
-Ramesh et al. [49] F - - - - $81.6$
-SSL4EVA [24] - - - - $86.9$
-EndoViT [5] 89.4 - - - - - - - - - - - - -
-SurgFM (ours) $87.0\pm 10.3$ $87.3\pm 7.5$ $87.7\pm 4.3$ $75.7\pm 9.2$ $83.1\pm 11.0$ $81.7\pm 10.9$ $\mathbf{86.3\pm 8.9}$ $69.8\pm 30.9$ $60.2\pm 26.1$ $82.6\pm 10.7$ $85.5\pm 9.9$ $87.3\pm 5.9$ $73.2\pm 9.4$ $66.8$
-SurgFM (ours) $\mathbf{92.7\pm 8.3}$ $\mathbf{93.0\pm 4.5}$ $\mathbf{93.1\pm 4.7}$ $\mathbf{85.1\pm 7.9}$ $\mathbf{90.6\pm 6.7}$ $\mathbf{85.2\pm 9.9}$ $82.9\pm 17.3$ $\mathbf{73.2\pm 32.0}$ $\mathbf{64.8\pm 27.3}$ $\mathbf{89.9\pm 7.1}$ $\mathbf{89.6\pm 6.5}$ $\mathbf{89.9\pm 6.9}$ $\mathbf{79.4\pm 9.9}$ $75.9$
+We cross-referenced the video titles with the predefined list of procedures.
+When exact title matches were not found, we leveraged the capabilities of the ChatGPTv4 API to perform a more nuanced analysis, incorporating a customized prompt (see Fig.
+2 in supplementary material).
+All annotations including the procedure type and the surgery type (robotic or non-robotic) were manually reviewed and validated to ensure the accuracy of the labels assigned.
+Table 2: Linear probing results for surgical phase recognition, surgical tool presence detection, and surgical action recognition tasks.
+The experiments are conducted with freezing the backbone and only fine-tuning a linear classification head to demonstrate the foundation model’s effectiveness.
+The predictions are computed on a frame-by-frame basis for all tasks.
+Method Pretraining data Phase recognition Tool detection Action recognition
+AutoLaparo Cholec80 M2CAI16 Cholec80 GraSP CholecT50
+Acc/F1 Acc/F1 Acc/F1 mAP mAP mAP
+MAE [21] LEMON 35.5/32.035.5/32.035.5 / 32.0 54.9/43.454.9/43.454.9 / 43.4 40.7/38.140.7/38.140.7 / 38.1 30.630.630.6 45.245.245.2 24.824.824.8
+Endo-FM [53] Private data 51.5/43.151.5/43.151.5 / 43.1 62.7/53.962.7/53.962.7 / 53.9 53.8/49.853.8/49.853.8 / 49.8 50.050.050.0 64.364.364.3 26.226.226.2
+EndoViT [5] Merged public data 45.4/38.745.4/38.745.4 / 38.7 46.9/39.046.9/39.046.9 / 39.0 38.5/37.238.5/37.238.5 / 37.2 34.734.734.7 47.147.147.1 29.529.529.5
+GSViT [48] GenSurgery 22.0/17.522.0/17.522.0 / 17.5 20.2/13.520.2/13.520.2 / 13.5 22.4/18.022.4/18.022.4 / 18.0 18.218.218.2 35.635.635.6 16.516.516.5
+SurgeNetXL [25] SurgeNetXL 68.8/57.068.8/57.068.8 / 57.0 73.2/65.173.2/65.173.2 / 65.1 67.8/61.967.8/61.967.8 / 61.9 72.172.172.1 62.762.762.7 45.345.345.3
+LemonFM LEMON $\mathbf{76.4/66.9}$ $\mathbf{75.8/68.6}$ $\mathbf{68.4/62.4}$ $\mathbf{76.8}$ $\mathbf{76.4}$ $\mathbf{50.4}$
+Figure 4: Proposed augmented distillation method.
+To encourage invariance in LemonFM to minor surgical motion and subtle appearance changes across patients, we introduce $W_{i}$ (Eq.
+1).
+The key component of $W_{i}$ is a pair of images (c) that are randomly selected from an augmentation pool (b).
+To populate the augmentation pool, which has capacity for four images, we first retrieve the nearest neighbors(1) of the input image from other videos of the same procedure type, but only include them in the pool if the cosine distance to the input image is smaller than 3×3$\times$3 × the distance between the input image and its preceding frame in the video.
+When not enough suitable neighbors are found, we supplement the pool with adjacent video frames(2)(3).
+3.2 Ground truth and proposed new downstream tasks
+Each of the 4194 surgical videos is annotated for two tasks: 1) multi-label procedure type classification, where the objective is to identify the procedure[s] performed in an intervention’s video from among 35 possible options; and 2) binary surgery type classification, where the goal is to indicate whether a surgery is conducted robotically or non-robotically.
+The rationale for these annotations is two-fold.
+Firstly, it allows us to validate the comprehensiveness of our dataset by ensuring that it encompasses a broad spectrum of surgical procedures.
+Secondly, it facilitates the development of new methods that use this information as a key input (e.g., our LemonFM foundation model, video classification/retrieval, video similarity measurement, and procedure-specific estimation of surgical remaining time).
+4 Proposed LemonFM Foundation Model
+In this section, we describe how we build LemonFM, our foundation model pretrained on LEMON using a novel augmented distillation approach.
+Following the standard practice in the field of surgical computer vision [51, 52, 43], we extract frames from the 4194 videos in LEMON at 1 fps for our foundation model pretraining.
 Proposed augmented distillation method.
-Inspired by DINO [8], we train a student network $f_{\boldsymbol{\theta_{s}}}$ to solve for the mapping $\sigma(f_{\boldsymbol{\theta_{t}}}(\boldsymbol{x}))=\sigma(f_{\boldsymbol{\theta_{s}}}(\boldsymbol{x}))$ where $\boldsymbol{x}\in\mathbb{R}^{H\times W\times 3}$ is an input image, $\sigma$ is the softmax function, $f_{\boldsymbol{\theta_{t}}}(\boldsymbol{x})\in\mathbb{R}^{C}$ is a target vector of dimension $C=2^{16}$ estimated from a teacher network $f_{\boldsymbol{\theta_{t}}}$, and $\boldsymbol{\theta_{s}}$ and $\boldsymbol{\theta_{t}}$ are the parameters of the student and teacher models, respectively.
+A foundation model for surgery should learn invariance to 1) minor movements of instruments and tissues observed between adjacent video frames, and 2) minimal appearance changes across patients undergoing the same surgical procedure (e.g., organ color variations).
+This is to encourage the model to produce similar embeddings for frames containing the same surgical tools and exhibiting comparable tissue appearances.
+Inspired by self-supervised knowledge distillation [8], we introduce a novel augmented distillation approach that leverages LEMON and employs ConvNeXt-L [36] for both the teacher and student networks.
+Specifically, we train a student network $f_{\boldsymbol{\theta_{s}}}$ to solve for the mapping $\sigma(f_{\boldsymbol{\theta_{t}}}(\boldsymbol{x}))=\sigma(f_{\boldsymbol{\theta_{s}}}(\boldsymbol{x}))$ where $\boldsymbol{x}\in\mathbb{R}^{H\times W\times 3}$ is an input image, $\sigma$ is the softmax function, $f_{\boldsymbol{\theta_{t}}}(\boldsymbol{x})\in\mathbb{R}^{C}$ is a target vector of dimension $C=2^{16}$ estimated from a teacher network $f_{\boldsymbol{\theta_{t}}}$.
+The student’s parameters $\boldsymbol{\theta_{s}}$ are updated by gradient descent, while the teacher’s parameters $\boldsymbol{\theta_{t}}$ are updated via exponential moving average from the student.
 We use the outputs of both networks to build two categorical distributions, $P_{s}(z|\sigma(f_{\boldsymbol{\theta_{s}}}(\boldsymbol{x})))=\sigma_{z}(f_{\boldsymbol{\theta_{s}}}(\boldsymbol{x}))$ and $P_{t}(z|\sigma(f_{\boldsymbol{\theta_{t}}}(\boldsymbol{x})))=\sigma_{z}(f_{\boldsymbol{\theta_{t}}}(\boldsymbol{x}))$, where $\sigma_{z}$ denotes the $z^{th}$ element of the softmax output, and $z\in\{1,\dots,C\}$.
 In training, we minimize the negative log-likelihood of $P_{s}(z|\sigma(f_{\boldsymbol{\theta_{s}}}(\boldsymbol{x})))$ following
-$\small\min_{\boldsymbol{\theta_{s}}}\sum_{i=1}^{I}$ $\sum_{\begin{subarray}{c}\boldsymbol{u}\in U_{i}\\ \boldsymbol{v}\in V_{i}\cup W_{i}\\ \boldsymbol{u}\neq\boldsymbol{v}\end{subarray}}\Big{[}d(\boldsymbol{u},\boldsymbol{v})-n(\boldsymbol{u},\boldsymbol{v})\Big{]}$ (1)
+$\min_{\boldsymbol{\theta_{s}}}\sum_{i=1}^{I}$ $\sum_{\begin{subarray}{c}\boldsymbol{u}\in U_{i}\\ \boldsymbol{v}\in V_{i}\cup W_{i}\\ \boldsymbol{u}\neq\boldsymbol{v}\end{subarray}}\Big{[}d(\boldsymbol{u},\boldsymbol{v})-n(\boldsymbol{u},\boldsymbol{v})\Big{]}$ (1)
 where $d(\boldsymbol{u},\boldsymbol{v})=\log\sum_{k^{\prime}=0}^{C-1}\exp\left(\sigma_{k^{\prime}}(f_{\boldsymbol{\theta_{s}}}(\boldsymbol{v}))\right)$, $n(\boldsymbol{u},\boldsymbol{v})=\sigma_{\arg\max_{z}P_{t}(z|\sigma(f_{\boldsymbol{\theta_{t}}}(\boldsymbol{u})))}(f_{\boldsymbol{\theta_{s}}}(\boldsymbol{v}))$, $I$ is the number of images in the dataset, $U_{i}$ contains two global augmentations of $\boldsymbol{x_{i}}$, and $V_{i}$ contains two global augmentations and four augmented crops of $\boldsymbol{x_{i}}$.
-A foundation model for surgery should learn invariance to 1) minor movements of instruments and tissues observed between adjacent video frames, and 2) minimal appearance changes across patients undergoing the same surgical procedure (e.g., organ color variations).
-Diverging from DINO  [8], and relying on the annotations available in Surg-3M, we introduce $W_{i}$ to provide a supervisory signal that encourages the model to learn these invariances.
-$W_{i}$ contains two images, and four augmented crops of these two images.
+Additionally, we introduce $W_{i}$ as an extra supervisory signal which contains two images, and four augmented crops of these two images.
 The procedure for selecting these two images is depicted in Fig.
 4.
-We use the $k$-nearest neighbors (in embedded space) of the input image from other videos of the same procedure type to learn invariance to appearance changes across patients, and the prior and subsequent video frames of the input image to learn invariance to the motion occurring between consecutive frames.
-We evaluate the performance of models trained with and without our augmented distillation (see Table 2) on surgical phase recognition (Cholec80), tool presence detection (Cholec80), and surgical action recognition (CholecT50).
-With augmented distillation, F1-score improved by $>1pp$ in both tool detection and action recognition tasks.
-Following these ablation results, we incorporated augmented distillation in our final SurgFM model.
-Figure 4: Proposed augmented distillation method.
-To encourage invariance in SurgFM to minor surgical motion and subtle appearance changes across surgeries, we introduce $W_{i}$ (Eq.
-1).
-The key component of $W_{i}$ is a pair of images (c) that are randomly selected from an augmentation pool (b).
-To populate the augmentation pool, which has capacity for four images, we first retrieve the nearest neighbors(1) of the input image from other videos of the same procedure type, but only include them in the pool if the cosine distance to the input image is smaller than 3×3$\times$3 × the distance between the input image and its preceding frame in the video that it was extracted from.
-When not enough suitable neighbors are found, we supplement the pool with adjacent video frames(2)(3).
+We use the $k$-nearest neighbors (in embedding space) of the input image from other videos of the same procedure type to learn invariance to appearance changes across patients, and the adjacent video frames of the input image to learn invariance to the motion occurring between consecutive frames.
+Table 3: Full fine-tuning results for surgical phase recognition, surgical tool presence detection, and surgical action recognition tasks.
+All experiments are conducted in a full fine-tuning setting (i.e., both backbone and head are fine-tuned).
+Type S denotes a specialist model trained for a specific task, whereas F denotes a foundation model whose generalization is assessed across multiple surgical downstream tasks.
+The superscript indicates the use of the standard 10-second relaxed boundary in phase transitions [51].
+We also report the performance of LemonFM when trained on only 50% of the labeled training data for the downstream tasks.
+Method Type Phase recognition Tool detection Action recognition
+AutoLaparo Cholec80 M2CAI16 Cholec80 GraSP CholecT50
+Acc/Jacc Acc/Jacc Acc/Jacc mAP mAP mAP
+TeCNO [12] S 77.3/50.777.3/50.777.3 / 50.7 88.6/75.188.6/75.188.6 / 75.1 86.1/74.486.1/74.486.1 / 74.4 -- -- --
+Trans-SVNet [30] 78.3/50.778.3/50.778.3 / 50.7 90.3/79.390.3/79.390.3 / 79.3 87.2/74.787.2/74.787.2 / 74.7 -- -- --
+LoViT [35] 81.4/55.981.4/55.981.4 / 55.9 92.4/81.292.4/81.292.4 / 81.2 /-/-- / - -- -- --
+MAE [21] F 63.5/42.263.5/42.263.5 / 42.2 84.6/71.284.6/71.284.6 / 71.2 77.8/63.677.8/63.677.8 / 63.6 72.372.372.3 53.053.053.0 33.133.133.1
+Endo-FM [53] 80.3/55.380.3/55.380.3 / 55.3 86.9/76.786.9/76.786.9 / 76.7 79.1/69.979.1/69.979.1 / 69.9 88.488.488.4 84.284.284.2 54.754.754.7
+EndoViT [5] 76.1/54.276.1/54.276.1 / 54.2 86.7/75.386.7/75.386.7 / 75.3 78.5/70.078.5/70.078.5 / 70.0 83.983.983.9 77.477.477.4 52.252.252.2
+GSViT [48] 60.5/37.060.5/37.060.5 / 37.0 75.3/57.875.3/57.875.3 / 57.8 55.8/39.655.8/39.655.8 / 39.6 34.334.334.3 37.737.737.7 26.326.326.3
+SurgeNetXL [25] 80.0/55.180.0/55.180.0 / 55.1 84.4/72.084.4/72.084.4 / 72.0 68.5/58.868.5/58.868.5 / 58.8 86.586.586.5 83.883.883.8 57.557.557.5
+LemonFM (ours) - 50% shot 84.9/62.484.9/62.484.9 / 62.4 92.1/78.792.1/78.792.1 / 78.7 86.1/77.786.1/77.786.1 / 77.7 91.591.591.5 89.489.489.4 60.160.160.1
+LemonFM (ours) $\mathbf{85.5/64.8}$ $\mathbf{92.7/85.1}$ $\mathbf{89.9/79.4}$ $\mathbf{93.7}$ $\mathbf{94.4}$ $\mathbf{61.9}$
 Proposed video classification model.
 Considering that each surgical procedure is localized to a specific region of the body, we posit that a characteristic scene can be associated with each procedure.
-Based on this assumption, and inspired by active learning [21], we propose SurgFM-Vid, a video classification method that employs SurgFM embeddings for video classification by aggregating them following
-$\small\boldsymbol{v_{e}}=\sum_{j=0}^{J-1}\omega_{j}\boldsymbol{\phi}_{j}$ (2)
+Based on this assumption, and inspired by active learning [19], we propose LemonFM-Vid.
+This is a video classification head that employs the frame embeddings produced by LemonFM to solve our proposed video classification downstream tasks.
+It does so by aggregating each frame’s embedding following
+$\boldsymbol{v_{e}}=\sum_{j=0}^{J-1}\omega_{j}\boldsymbol{\phi}_{j}$ (2)
 where $\boldsymbol{v_{e}}$ is a video embedding, $J$ is the number of frames in the video, $\omega_{j}=\frac{t(\boldsymbol{\phi}_{j})}{\epsilon+\sum_{m=0}^{J-1}t(\boldsymbol{\phi}_{m})}$, $\boldsymbol{\phi}_{j}$ is the embedding of frame $j$, $t(\boldsymbol{\phi})=\left(\frac{1}{K}\sum_{\boldsymbol{\eta}\in K-NN(\boldsymbol{\phi})}d(\boldsymbol{\phi},\boldsymbol{\eta})\right)^{-1}$ is the typicality [19] of a frame, $K-NN(\boldsymbol{\phi})$ with $K=20$ are the nearest neighbors of $\boldsymbol{\phi}$, $d$ is the cosine distance, and $\epsilon=10^{-8}$.
 Then, each $\boldsymbol{v_{e}}$ is classified with a single-layer MLP.
-5 Evaluation
-5.1 Downstream Tasks, Datasets and Metrics
-We evaluate SurgFM on three surgical downstream tasks: surgical phase recognition [54, 53, 57, 4], surgical tool presence detection [54, 4], and surgical action recognition [48, 4].
+5 Experiments
+5.1 Downstream Tasks, Datasets and Evaluation
+We evaluate LemonFM on four surgical downstream tasks: surgical phase recognition [52, 51, 50], surgical tool presence detection [51, 4], surgical action recognition [43], and surgical semantic segmentation [23].
 To do so, we use the most widely adopted datasets, data splits, and evaluation metrics.
 Surgical phase recognition.
-We use M2CAI16 [54], Cholec80 [54], AutoLaparo [57], and GraSP [4].
-For this task, we report accuracy, precision, recall, Jaccard, F1-score, and mean Average Precision (mAP).
-The accuracy for a video is computed by dividing the number of frames whose class has been correctly predicted by the total number of frames.
-The overall accuracy for a dataset is defined as the mean accuracy value over all videos, henceforth referred to as video-level accuracy [28, 29, 35, 27].
-Precision, recall, and Jaccard are computed differently than accuracy [28, 29, 35, 27].
-These three phase-level metrics are computed for each class and video, then averaged: first within videos, and secondly across classes.
-The F1-score, computed for Cholec80, is also computed for each class and video, but averaged first over classes and then over videos, and referred to as video-level F1-score [49, 24].
-For GraSP, the frame-level mAP was calculated by averaging the mAP scores across all frames in the test set [4].
-Surgical tool presence detection. Cholec80 and GraSP are utilized in the surgical phase recognition task. We report mAP across all frames in the testing dataset [49, 4].
+For this task, we use three datasets: AutoLaparo [52], M2CAI16 [51], and Cholec80 [51].
+These datasets include complete surgical workflows with frame-level phase annotations, where each frame corresponds to a single phase.
+We perform experiments with both linear probing, where we freeze the backbone and only fine-tune a linear classification head, and full fine-tuning, where both the backbone and the head are fine-tuned.
+We report the frame-wise accuracy and F1-score for the linear probing setting (Table 2), and the video-level accuracy and Jaccard (more details in supplementary material) for the full fine-tuning setting.
+To produce video-level online predictions, we replace the linear head with an MS-TCN head [12] (Table 3).
+Surgical tool presence detection.
+Cholec80 and GraSP [4] are utilized in the surgical tool presence detection task.
+These datasets contain surgical tool presence labels for each frame, allowing multiple tools to be present in a single frame.
+We report mean Average Precision (mAP) across all frames for both the linear probing and full fine-tuning settings (see Table 2 and 3).
+We use a linear head for both experiments.
 Surgical action recognition.
-For this task we use CholecT50 [44] and SAR-RARP50 [48].
-For CholecT50, we report the video-level mAP, which is computed by first calculating the average mAP for each class across all videos and then taking the overall mean of these class averages [44].
-For SAR-RARP50 [48], we report video-level accuracy and video-level F1@10 [48].
-5.2 Downstream Task Evaluation Protocols
-To capture temporal correlations in the surgical phase recognition downstream task, we couple our SurgFM backbone with a TCN head [12].
-This TCN head takes both current and previous frames into account, but excludes future frames.
-For a fair comparison with previous methods [26, 54, 30, 12, 27, 35], we run our evaluation in Cholec80 and M2CAI16 with the standard 10-second relaxed boundary between some of the surgical phases [54].
-Also to be able to compare with previous literature [57, 35, 4], we do not use a relaxed boundary for AutoLaparo and GraSP.
-For surgical tool presence detection, we use a linear head for all datasets.
-For surgical action recognition we use a different head for each dataset.
-To evaluate on CholecT50, we utilize a linear head.
-The predictions are done per-frame, no temporal information is used, as in the approaches that we benchmark against [44, 43].
-For SAR-RARP50, we evaluate our model in two different modes: online, where only previous frames are available (TCN head [12]), and offline ($\pm$ subscript in Table 5), which incorporates information from both previous and future frames (MS-TCN head [33]), to ensure a fair comparison with all competing methods in [48].
-In our evaluation, we compare three types of models: generalist (G), pretrained on ImageNet and fine-tuned for each downstream task; specialist (S), designed for specific downstream tasks; and foundational (F), pretrained in a self-supervised manner and fine-tuned on each downstream task.
-Table 4: Performance evaluation on surgical tool presence detection.
-We evaluate the performance on two datasets: Cholec80 [54], and GraSP [4].
-The metrics are computed as described in Sec.
-5.1.
-The evaluation protocol and the meanings of G, S, and F in the type column are described in Sec.
-5.2.
-The superscript denotes that the backbone is frozen during fine-tuning.
-Method Type Cholec80 GraSP
-mAP (%) mAP (%)
-ConvNeXt-L [36] G 69.3 $68.1$
-ConvNeXt-L [36] $93.3$ $89.5$
-EndoNet [54] S $81.0$
-MTRCNet-CL [27] $89.1$
-Rendezvous [44] - $80.4$
-TAPIS [4] - $\mathbf{95.1}$
-Ramesh et al. [49] F $93.6$
-SurgFM (ours) $76.1$ $76.4$
-SurgFM (ours) $\mathbf{93.7}$ $94.4$
-6 Results and Discussion
+We use CholecT50 [43] as the downstream dataset.
+Surgical action labels are annotated at frame level, enabling multiple actions occur simultaneously.
+We use a linear head for both the linear probing and full fine-tuning settings and apply the same metrics as in the surgical tool presence detection task.
+Surgical semantic segmentation.
+We employ the CholecSeg8k [23] dataset, which contains pixel-wise semantic annotations for each frame.
+The experiments are conducted in full fine-tuning using a feature pyramid network decoder [32].
+Dice score is reported over all frames in the dataset (see Table 4).
+Table 4: Evaluation on surgical semantic segmentation.
+The experiments are conducted with using a feature pyramid network (FPN) decoder [32] in a full fine-tuning manner.
+The superscript denotes the LemonFM backbone is frozen.
+We also report the performance of LemonFM when trained on only 50% of the labeled training data.
+Method Pretraining data CholecSeg8k
+mDice
+MAE LEMON 68.568.568.5
+Endo-FM Private data 54.554.554.5
+EndoViT Merged public data 73.073.073.0
+GSViT GenSurgery 53.453.453.4
+SurgeNetXL SurgeNetXL 70.070.070.0
+LemonFM LEMON 71.971.971.9
+LemonFM - 50% shot LEMON 78.278.278.2
+LemonFM LEMON $\mathbf{81.3}$
+5.2 Comparison with SotA
+The performance of our LemonFM compared to the state of the art in each downstream task and dataset is shown in Fig.
+1, and Tables 2, 3 and 4.
+These results demonstrate that the diversity of procedures in LEMON enables LemonFM to effectively extract features that are useful for different surgical procedures and downstream tasks.
+Surgical phase recognition.
+As shown in Table 3, LemonFM outperformed every other surigcal foundation model in AutoLaparo ($>$9pp Jacc.), Cholec80 ($>$8pp Jacc.), and M2CAI16 ($>$9pp Jacc.).
+Furthermore, it also surpassed state-of-the-art specialist models for these datasets.
+The advantage of LemonFM persists under the linear probing setting (Table 2), showing that the pretrained features are inherently informative and transfer well without additional fine‑tuning.
+Surgical tool presence detection. As shown in Table 2 and 3, LemonFM outperformed all surgical foundation models in Cholec80 and GraSP.
+Surgical action recognition. LemonFM outperformed all surgical foundation models in CholecT50 (Table 2 and 3).
+Surgical semantic segmentation. As shown in Table 4, LemonFM surpassed every other foundation model in the pixel-wise semantic segmentation task by a large margin ($>$8pp mDice).
+Low data regime. As shown in Tables 3 and 4, when trained on 50% of the data, LemonFM outperforms other surgical foundation models fine-tuned on 100% of the data, demonstrating high data efficiency.
 5.3 Ablation Studies
-SurgFM design choices.
-We performed incremental ablation studies that test the choice of backbone (ViT vs ConvNeXt), self-supervision method (DINO vs DINOv2 vs MAE), and our proposed augmented distillation (Sec.
-4).
-The results are shown in Table 2.
-DINO outperformed DINOv2 and MAE in all the downstream tasks and metrics evaluated, with improvements $>$8 percentage points (pp) in mAP in both tool detection and action recognition.
-ConvNeXt-L achieved comparable results to ViT-L in phase recognition, and improved $>$2pp (F1-score) in surgical tool detection and $>$1pp (mAP) in surgical action recognition.
-Our proposed augmentation achieved comparable results in phase recognition, and improved $>$1pp (mAP) in tool detection and $>$1pp (F1-score) in action recognition.
-Pretrained weights.
-We compare the effects of employing pretrained weights from ImageNet in the ConvNeXt-L backbone with those of using Surg-3M.
-As shown in Tables 3, 4, and 5, SurgFM (which also has a ConvNeXt-L backbone, but is pretrained on Surg-3M) outperformed the ConvNeXt-L backbone pretrained on ImageNet in all downstream tasks and datasets.
-This finding suggests that SurgFM serves as a superior starting point for surgery-specific downstream task fine-tuning compared to models pretrained on general-purpose datasets like ImageNet.
-This is clearly shown in the surgical tool presence detection experiments (Table 4) on the GraSP dataset, where SurgFM outperformed a ConvNeXt-L pretrained on ImageNet by more than 4pp in mAP.
-6.2 Comparison with SotA
-The performance of our SurgFM compared to the state of the art in each downstream task and dataset is shown in Fig. 1 and Tables 3, 4, and 5.
-Table 5: Performance evaluation on surgical action recognition.
-Two datasets are used: CholecT50 [42] and SAR-RARP50 [48].
-The metrics are computed as described in Sec.
-5.1.
-The evaluation protocol and the meanings of G, S, and F are explained in Sec.
-5.2.
-The superscript denotes that the SurgFM backbone is frozen during fine-tuning.
-The subscript ± indicates that both past and future frames are used [48].
-Method Type CholecT50 SAR-RARP50
-mAP (%) Accuracy (%) F1@10 (%)
-ConvNeXt-L± [36] G $62.9\pm 2.3$ $71.9$ $70.4$
-Tripnet [43] S 58.8$\pm$3.1 - -
-Rendezvous [44] $60.4\pm 2.8$
-Attention Tripnet [44] $61.1\pm 2.0$
-SummerLab-AI± [48] - $\mathbf{81.5}$ $\mathbf{84.1}$
-Uniandes± [48] - $78.6$ $82.3$
-MS-TCN± [48, 16] - $69.0$ $70.7$
-LoViT± [48, 35] - $59.8$ $43.0$
-SurgFM (ours) F $\mathbf{64.2\pm 1.7}$ $70.2$ $68.4$
-${}_{\pm}^{*}$ (ours) $63.7\pm 1.7$ $72.4$ $74.3$
-SurgFM± (ours) $\mathbf{64.2\pm 1.7}$ $74.1$ $74.3$
-Surgical phase recognition.
-As shown in Table 3, SurgFM outperformed all specialist [54, 27, 52, 26, 62, 12, 20, 29, 30, 35] and foundation [49, 24, 5] models in Cholec80 ($>$3pp Jacc.), AutoLaparo ($>$8pp Jacc.), and M2CAI16 ($>$4pp Jacc.).
-These results demonstrate that SurgFM is able to effectively extract features that are useful for different surgical procedures, capitalizing on the diversity of surgery types present in the Surg-3M dataset.
-In GraSP, SurgFM demonstrated performance comparable (with an mAP difference of $<$1pp) to that of the methods presented in [17, 55, 4].
-However, these competing methods [17, 55, 4] utilize both past and future frames (offline), whereas our SurgFM is restricted to observing only the current and preceding frames, reflecting its real-time operational focus, and its broader range of application [35].
-Despite being constrained by its limited temporal perspective, SurgFM achieves comparable performance with these more information-rich approaches.
-Surgical tool presence detection.
-As shown in Table 4, SurgFM outperformed all specialist [54, 27, 44] and foundation [49] models in Cholec80, and achieved a comparable performance to [44, 4] in GraSP.
-Surgical action recognition.
-As shown in Table 5, SurgFM outperformed all specialist model [43, 44] in CholecT50 ($>$3pp mAP), and achieved a competitive performance to [48] in SAR-RARP50 ($<$7pp difference in accuracy).
-Table 6: Leaderboard for the proposed tasks on Surg-3M. Multi-label (35 classes) video classification of procedure types and binary classification of surgery types.
+We conducted incremental ablation studies on the surgical phase recognition (AutoLaparo) and surgical semantic segmentation (CholecSeg8k) tasks (see Table 5).
+Table 5: Ablation studies.
+We froze the backbone and only fine-tuned the task-specific heads (i.e., a linear head for phase recognition and a FPN head for semantic segmentation).
+The column “Curation” marks whether the raw surgical videos were processed with the pipeline in Fig.
+2.
+The column “Aug” shows whether the augmented distillation method in Fig.
+4 was applied.
+Dataset Curation Aug Backbone AutoLaparo CholecSeg8k
+Acc/F1 mDice
+IN-1K N/A ConvNeXt-L 63.6/53.063.6/53.063.6 / 53.0 64.464.464.4
+LEMON ConvNeXt-L 71.7/61.471.7/61.471.7 / 61.4 67.467.467.4
+LEMON ConvNeXt-L 75.3/65.975.3/65.975.3 / 65.9 68.768.768.7
+LEMON ViT-L 75.6/66.175.6/66.175.6 / 66.1 61.261.261.2
+LEMON ConvNext-B 74.7/64.574.7/64.574.7 / 64.5 68.568.568.5
+LEMON ConvNeXt-L $\mathbf{76.4/66.9}$ $\mathbf{71.9}$
+The significance of LEMON.
+We compare the effects of using LEMON for pretraining the surgical foundation model against those obtained with ImageNet-1K [13] (row 1 and row 3 in Table 5).
+The model trained on LEMON outperformed the one pretrained on ImageNet by a large margin (+12.9pp F1-score on phase recognition, +4.3pp mDice on semantic segmentation).
+This finding suggests that pretraining on LEMON offers a superior approach for surgery-specific foundation model training compared to pretraining on general-purpose datasets like ImageNet.
+The importance of rigorous data curation.
+We analyze the importance of our data curation pipeline by comparing LemonFM against a model pretrained on an uncurated version of LEMON (i.e., the raw surgical YouTube videos without our data curation).
+Our data curation pipeline (Fig.
+2) yields substantial improvements.
+Specifically, the curated dataset leads to an increase of 4.5pp F1-score in the phase recognition task and 1.3pp of mDice in the semantic segmentation task (rows 2 and 3 in Table 5).
+The effectiveness of the augmented distillation.
+We compared the effectiveness of employing our proposed augmented distillation method (Fig.
+4) against a vanilla DINO [8].
+LemonFM consistently outperformed the vanilla DINO baseline, with the largest gain (+3.2pp mDice) in semantic segmentation (rows 3 and 6 in Table 5).
+LemonFM design choices.
+We performed ablation studies to evaluate the effects of two design choices for surgical foundation model pretraining: backbone architecture, ViT vs.
+ConvNeXt, and self-supervision method, generative-based [21] vs.
+discriminative-based [8].
+As shown in row 4 and 6 of Table 5, when pretrained on LEMON with our augmented knowledge distillation, the ConvNeXt backbone consistently outperformed the ViT-based counterpart.
+The improvement in performance is particularly notable in the pixel-wise semantic segmentation task (+10.7pp mDice).
+In Tables 2, 3, and 4, we compare methods whose pretext task is generative-based (MAE and EndoViT) against methods whose pretext task is discriminative-based (classification) such as Endo-FM, SurgeNetXL and our LemonFM.
+The results show that discriminative‑based models yield a substantial performance advantage, especially when the backbone is frozen, indicating generative objectives such as MAE are less effective at capturing surgical visual cues.
+Table 6: Leaderboard for the proposed tasks on LEMON. Multi-label (35 classes) video classification of procedure types and binary classification of surgery types.
 Method Procedure type Surgery type
-mAP (%) F1 (%) Accuracy (%) F1 (%)
-SlowFast [17] 22.0 23.9 88.5 87.5
-TimeSformer [6] 42.1 37.5 93.2 92.7
-MViTv2 [34] 49.5 41.8 95.8 94.6
-Video Swin Transformer [37] 51.4 47.9 98.8 98.7
-SurgFM-Vid (ours) $\mathbf{57.8}$ $\mathbf{49.3}$ $\mathbf{98.9}$ $\mathbf{98.9}$
-Frozen SurgFM model.
-As shown in Tables 3, 4, and 5, the frozen version of SurgFM, denoted SurgFM, is indeed a foundation model.
-It works on all datasets and tasks without retraining.
-In Cholec80, SurgFM outperforms specialist SotA architectures such as EndoNet, PhaseNet, and SV-RCNet, OHFM, TeCNO.
-In AutoLaparo, SurgFM outperforms all the six specialist models we compared it with.
-In M2CAI16, SurgFM outperforms two specialist SotA models, PhaseNet and SV-RCNet.
-In SAR-RARP50, SurgFM outperforms MS-TCN and LoViT.
-Data scaling.
-We investigated the extent to which annotation effort can be alleviated by leveraging our pretrained SurgFM model.
-To do so, we fine-tuned it on downstream tasks with 10%, 50%, and 100% of the annotated data (Fig.
-5).
-With 50% of the data, the performance decreased a maximum of 4.4pp across all datasets and tasks except GraSP, which has a total of four videos for training at 100%, and where performance dropped by 9.6pp.
-The closest result was achieved in phase recognition (Cholec80), with SurgFM achieving an accuracy of 92.3% on 50% of the data, only 0.1pp below prior SotA with 100% of the data [35].
-Also at 50%, our SurgFM outperformed prior SotA [35] in AutoLaparo at 100%.
-These results suggest that the features learned by SurgFM from Surg-3M are highly informative.
-Comparing our results with other foundation models reported in the literature, SurgFM achieved $>$6pp (F1-score) than Ramesh et al.
-[49] in phase recognition (Cholec80), despite using only 10% of the data compared to their 12.5%.
-Furthermore, SurgFM also outperformed Ramesh et al.
-by $>$6pp (mAP) on tool presence detection (Cholec80) at 10%.
+mAP/F1 Acc/F1
+SlowFast 22.0 /23.9 88.5 /87.5
+TimeSformer 42.1/37.5 93.2 /92.7
+MViTv2 49.5/41.8 95.8 /94.6
+Video Swin Transformer 51.4/47.9 98.8/98.7
+LemonFM-Vid $\mathbf{57.8/49.3}$ $\mathbf{98.9/98.9}$
 5.4 Leaderboard
-To establish a baseline for the novel tasks presented in Sec.
-3, we conducted an evaluation of SotA approaches on Surg-3M (Table 6), which serves as a benchmark for future research endeavors.
-By leveraging the robust representations extracted from SurgFM, our video classification model, SurgFM-Vid (Sec.
-4), demonstrated significant improvements ($>$6 mAP) over all the leading video classification models tested on the multi-label procedure type classification task.
-On the binary surgery type classification task, our approach achieved comparable accuracy to the top performer, the Video Swin Transformer [37].
-The incorporation of frame-level features extracted by SurgFM, in conjunction with the strategic application of prior knowledge through weighted feature embeddings informed by frame typicality, leads to substantial performance gains, underscoring the value of SurgFM as a critical component in surgical video analysis tasks.
-Despite these improvements, our results also highlight the challenges associated with surgical procedure classification, underscoring the need for continued innovation in this area.
+We benchmarked state-of-the-art video classification models on LEMON (Table 6), including SlowFast [15], TimeSformer [6], MViTv2 [34] and Video Swin Transformer [37].
+Leveraging the robust representations extracted from LemonFM, our video classification model, LemonFM-Vid (Sec.
+3), improved multi‑label procedure classification by 6.4pp mAP over all leading models and matched the top performance of the Video Swin Transformer on the binary surgery type classification task.
+The incorporation of frame-level features extracted by LemonFM, in conjunction with the weighted feature embeddings informed by frame typicality, led to substantial performance gains, underscoring the value of LemonFM for surgical video analysis tasks.
+Nonetheless, the mAP score of 57.8% achieved by LemonFM-Vid (the top performer method) highlights the challenges associated with our newly proposed procedure type video classification task, underscoring the need for continued innovation in this area.
 6 Conclusion
-Figure 5: Data scaling.
-Results of SurgFM when fine-tuned on the downstream tasks with different percentages of annotated training samples.
-The radius of the circles, computed following $2\log(\sigma+1)$, represents the standard deviation ($\sigma$) of each metric estimated over five cross-validation folds.
-Prior SotA is computed using 100% of the available data.
-Best viewed online.
-Building on a large corpus of publicly available surgical videos, this work has two main contributions: we have created a large-scale and diverse surgical video dataset, Surg-3M, and a foundation model, SurgFM.
-Our SurgFM outperformed specialist architectures on four datasets and three downstream tasks, demonstrating its potential as a valuable tool for computer vision research in surgery.
-Furthermore, our experimental results show that SurgFM can maintain its performance with significantly reduced data annotation needs (50% or even 10% for some tasks), highlighting the efficiency of our approach.
-This work showcases the value of Surg-3M and SurgFM as a foundation for future research in surgical computer vision.
-To further improve these contributions, we plan to address key limitations of our current approach, including adapting SurgFM to pixel-wise tasks and enhancing the self-distillation approach to better adapt to surgical data.
-Ultimately, this work highlights the importance of publicly available surgical videos in developing more effective models for computer vision research in surgery, and we believe that Surg-3M and SurgFM will serve as catalysts for future advancements in this field.
+Building on a large corpus of publicly available surgical videos, in this work we presented three main contributions: 1) we proposed a novel data curation pipeline to curate online surgical videos; 2) leveraging this pipeline, we created a large-scale and diverse surgical dataset, LEMON; and 3) we developed a foundation model specific for surgical applications, LemonFM.
+Our results show that LemonFM outperformed both state-of-the-art surgical foundation models and task-specific architectures across diverse surgical downstream tasks.
+Even when fine-tuned with only half of the labeled data, LemonFM maintained its advantages.
+These findings demonstrate LemonFM’s effectiveness and verify the quality of LEMON as a valuable resource for computer vision research in surgery.
+We expect that both LEMON and LemonFM will serve as a foundation for future research in surgical computer vision and facilitate the advancement of autonomous robotic surgery systems.
+Future work will concentrate on the development of a surgery-specific video foundation model.
 […]
 Supplementary Material
-You are a highly knowledgeable assistant specializing in surgical procedures and medical terminology.
-Your expertise includes identifying and categorizing surgical interventions based on clinical
-descriptions and procedural contexts.
-Here is a list of 35 possible surgical procedure types: pancreatectomy, pancreaticoduodenectomy,
-splenectomy, ampullectomy, hepatectomy, nephrectomy, low anterior resection, colectomy,
-abdominoperineal resection, pulmonary lobectomy, hartmanns, prostatectomy, gastric bypass,
-duodenal switch, gastrectomy, small bowel resection, hernia repair, ulcer repair, cholecystectomy,
-appendectomy, ileocolic resection, cecectomy, myomectomy, hysterectomy, nissen fundoplication,
-adrenalectomy, thymectomy, rectopexy, adhesiolysis, esophagectomy, cystectomy, jejunostomy, ileorectal
-anastomosis, kidney transplant, vaginectomy.
-Based on the description of the surgical video: <video title>, determine the most likely procedure type
-from the list. Focus on matching the description to the procedure type that best aligns with the
-terminology and context provided.
-Figure S1: ChatGPT prompt employed to match video titles to procedure types. The title of the video to be matched is inserted where the $<$video title$>$ tag is located.
+Figure 1: Diversity and procedure prevalence in LEMON.
+Representative samples from various procedures, demonstrating the diverse range of cases in our curated dataset (left, right).
+Distribution of surgical frames by procedure type (center).
 Appendix A Surgical Background
@@
 In contrast, non-robotic (also known as traditional or conventional) laparoscopy requires the surgeon to directly hold and steer the laparoscopic hand-held instruments (i.e., no robotic arms involved).
-Appendix B Details of Data Curation
-This section provides details of the implementation of our data curation pipeline.
+Appendix B Data Curation Details
+This section provides details of our data curation pipeline and presents LEMON’s procedure diversity and prevalence (Fig. 1).
 B.1 Video Classification
 Video summarization.
-To improve the efficiency of annotating surgical and non-surgical videos, we obtained 4$\times$4-image video storyboards (i.e.
-a single image containing a collage of key video frames) for all the collected videos using the method described in [18].
+To improve the efficiency of annotating surgical and non-surgical videos, we obtained 4$\times$4-image video storyboards (i.e., a single image containing a collage of key video frames) for all the collected videos using the method described in [16].
 Storyboards enabled us to quickly determine whether a video contained substantial surgical footage, thereby avoiding the need for complex analysis on the entire sequence.
@@
 To ensure the accuracy of the inference results, we manually reviewed all the videos classified as surgical.
-Performance of video storyboard classification models.
-We trained five video storyboard classification models (different data splits) to categorize the videos as either surgical or non-surgical.
-Each fold was split into training, validation, and testing sets with ratios of 0.8, 0.1, and 0.1, respectively.
-The average F1-score of the video storyboard classification models was $94.75\%\pm 1.1$.
-The results for each cross-validation fold are shown in Table S1.
-Table S1: Performance of the video storyboard classification models across five folds.
+[…]
+Table 1: Performance of the video storyboard classification models across five folds.
 Fold F1-score (%) Accuracy (%) Precision (%) Recall (%)
@@
 The resulting videos were manually quality controlled.
-Performance of frame classification models.
-We trained five frame classification models to classify a video frame as either surgical or non-surgical.
-Each model of the five was trained in a different training-validation-testing split of the data, with a split rate of 0.8, 0.1, and 0.1.
-As shown in Table S2, the average F1-score of the frame classification models was $95.64\%\pm 0.94$.
-Table S2: Performance of frame classification models across five folds.
+[…]
+B.3 Video Preprocessing
+We use our trained frame classifier to detect and remove the intraoperative non-surgical frames.
+Additionally, we manually annotated 2719 surgical frames with 4584 non-surgical bounding-box instances and trained a YOLOv8 Nano model [31] to detect and obliterate the non-surgical content in surgical frames.
+After this process, we manually quality control all the remaining curated videos.
+[…]
+You are a highly knowledgeable assistant specializing in surgical procedures and medical terminology.
+Your expertise includes identifying and categorizing surgical interventions based on clinical descriptions and procedural contexts.
+Here is a list of 35 possible surgical procedure types: pancreatectomy, pancreaticoduodenectomy, splenectomy, ampullectomy, hepatectomy, nephrectomy, low anterior resection, colectomy, abdominoperineal resection, pulmonary lobectomy, hartmanns, prostatectomy, gastric bypass, duodenal switch, gastrectomy, small bowel resection, hernia repair, ulcer repair, cholecystectomy, appendectomy, ileocolic resection, cecectomy, myomectomy, hysterectomy, nissen fundoplication, adrenalectomy, thymectomy, rectopexy, adhesiolysis, esophagectomy, cystectomy, jejunostomy, ileorectal anastomosis, kidney transplant, vaginectomy.
+Based on the description of the surgical video: <video title>, determine the most likely procedure type from the list.
+Focus on matching the description to the procedure type that best aligns with the terminology and context provided.
+Figure 2: ChatGPT prompt employed to match video titles to procedure types. The title of the video to be matched is inserted where the $<$video title$>$ tag is located.
+B.4 Video Annotation
+For the surgery type, a video is considered to be robotic if the video title includes any of the following keywords: Robotic, Robot, Robo, Hugo, Versius, Senhance, Telerobotic, Console, and da Vinci.
+The search for these terms in the video titles was case-insensitive.
+The remaining videos, with titles that do not include any of these keywords, are manually verified to ensure that they are manual surgical procedures.
+For the surgical procedure type, we cross-reference the video titles with the predefined list of procedures.
+When exact title matches are not found, we leverage the capabilities of the ChatGPTv4 API to perform a more nuanced analysis, incorporating a customized prompt as shown in Fig.
+2.
+All annotations are manually quality controlled.
+[…]
+Table 2: Performance of frame classification models across five folds.
 Fold F1-score (%) Accuracy (%) Precision (%) Recall (%)
@@
 Std Dev 0.94 0.88 0.49 1.39
-B.3 Video Annotation
-For the surgery type, a video is considered to be robotic if the video title includes any of the following keywords: Robotic, Robot, Robo, Hugo, Versius, Senhance, Telerobotic, Console, and da Vinci.
-The search for these terms in the video titles was case-insensitive.
-The remaining videos, with titles that do not include any of these keywords, are manually verified to ensure that they are manual surgical procedures.
-For the surgical procedure type, we cross-reference the video titles with the predefined list of procedures.
-When exact title matches are not found, we leverage the capabilities of the ChatGPTv4 API to perform a more nuanced analysis, incorporating a customized prompt as shown in Fig.
-S1.
-All annotations are manually quality controlled.
-After curation, the final Surg-3M dataset resulted in $4194$ surgical videos, 94% of which are $1280\times 720p$, the remaining $6\%$ have varied resolutions, with a minimum of $640\times 480$ pixels.
-Appendix C SurgFM Pretraining: Implementation Details
-This section provides an overview of the pretraining process for the SurgFM model, and the image preprocessing pipeline employed.
-SurgFM pretraining.
-For the pretraining of SurgFM, we utilized eight NVIDIA V100 GPUs (32 GB) with a batch size of 24 per GPU.
-We employed AdamW as the optimizer with a teacher temperature of 0.04, fp16 precision, an initial learning rate of 5e-4 (after warm-up), and a minimum learning rate of 1e-6.
+Appendix C LemonFM Pretraining Details
+This section provides an overview of the pretraining process for the LemonFM model.
+LemonFM pretraining.
+For LemonFM pretraining, we ran on an Ubuntu 22.04.5 LTS node with eight NVIDIA V100 GPUs (32GB each), using a batch size of 24 per GPU under PyTorch 2.5.1+cu124 (CUDA 12.4).
+We employed AdamW as the optimizer with a teacher temperature of 0.04, fp16 precision, an initial learning rate of 5e-4 (after warm-up), a minimum learning rate of 1e-6, and a random seed of 30.
 The model was trained for 60 epochs, including 10 warm-up epochs.
@@
 The model with the lowest training loss was selected as the final model.
-Figure S2: Image preprocessing in SurgFM. We automatically detect and crop out the user interface information from the surgical frames.
-Image preprocessing. We utilize our trained non-surgical bounding-box detection model to identify and obliterate UI elements, as demonstrated in Fig. S2.
-Performance of non-surgical content detection models.
-We trained five non-surgical content detection models to detect and obliterate non-surgical regions in surgical video frames.
-Each model was trained on a different training-validation-testing split of the data, with a split ratio of 0.8, 0.1, and 0.1, respectively.
-The average mAP50 of the five models was $79.29\%\pm 3.4$, and the average mAP50-95 was $66.18\%\pm 2.98$.
-The results for all the cross-validation folds are shown in Table S3.
-Table S3: Performance of bounding box detection models across five folds.
-Fold Images Instances Box Precision (%) Box Recall (%) mAP50 (%) mAP50-95 (%)
+Appendix D Downstream Task Details
+In this section, we provide details on the datasets used for evaluating each downstream task, including their respective data splits, as well as training configurations for linear probing and full fine-tuning settings.
+The experiments were conducted on an Ubuntu 22.04.5 LTS node equipped with an Intel IceLake Xeon CPU (72 vCPUs) and two NVIDIA A100 GPUs (each with 80 GB of memory), using PyTorch 2.5.1+cu124 (CUDA 12.4).
+D.1 Data splits
+Surgical phase recognition.
+For AutoLaparo, we followed the standard split of ten training videos, four for validation, and seven for testing [52].
+For M2CAI16, we followed the data splits in [26, 51, 30], dividing the dataset into 27 training videos and 14 testing videos, respectively.
+For Cholec80, we adopted the data splits in [51, 12, 35, 27], allocating 40 videos for training, eight for validation, and 32 for testing.
+Surgical tool presence detection.
+For GraSP, we adopted the data splits specified in [4], with four videos for training, four for validation, and five for testing.
+The data split for Cholec80 is identical to those employed in the surgical phase recognition task.
+Surgical action recognition. For CholecT50, we followed the formal data split as proposed in [43].
+Surgical semantic segmentation. For CholecSeg8k, we followed previous works [25, 18] and use 75% of the videos for training and 25% of the videos for testing (videos 12, 20, 48 and 55).
+D.2 Linear probing
+We used the teacher model from LemonFM as our backbone.
+The images were resized to $\times 224$ 224 for all downstream tasks.
+The downstream models with linear heads were trained with a batch size of 512, an initial learning rate (LR) of 1e-3, the AdamW optimizer, random seed of 30, and cross-entropy loss.
+An early stopping criterion was applied, stopping training after 10 epochs if the validation loss shows no improvement.
+Table 3: Performance of bounding box detection models across five folds.
+Fold Images Instances Precision Recall mAP50 mAP50-95
 Fold 0 272 447 74.79 77.63 75.50 62.00
@@
 Std Dev 4.41 2.33 3.40 2.98
-Appendix D Downstream Tasks: Data Splits and Fine-tuning
-In this section, we provide the details of the datasets used to evaluate each downstream task along with their corresponding data splits.
-Surgical phase recognition.
-For M2CAI16, we follow the data splits in [26, 54, 30], dividing the dataset into 27 training videos and 14 testing videos, respectively.
-For Cholec80, we adopt the data splits in [54, 12, 35, 27], allocating 40 videos for training, eight for validation, and 32 for testing.
-For AutoLaparo, we follow the standard split of ten training videos, four for validation, and seven for testing [57].
-For GraSP, we adopt the data splits specified in [4], with four videos for training, four for validation, and five for testing.
-Surgical tool presence detection. The data splits for Cholec80 and GraSP are identical to those employed in the surgical phase recognition task.
-Surgical action recognition.
-For CholecT50, we follow the standard 5-fold data splits as proposed in [42].
-For SAR-RARP50 [48] we use its recommended split of 27 videos for training, 13 for validation, and 10 for testing.
-Fine-tuning on Downstream tasks.
-We used the teacher model from SurgFM as our backbone and employed three different protocols for downstream task fine-tuning.
-The images were resized to $448\times 448$ for all downstream tasks.
-The downstream models with linear heads were trained with a batch size of 112, an initial learning rate (LR) of 1e-4, the AdamW optimizer, and cross-entropy loss.
+D.3 Full fine-tuning
+We used the teacher model from LemonFM as our backbone and employed three different protocols for downstream task fine-tuning.
+The images were resized to $\times 224$ 224 for all downstream tasks.
+The downstream models with linear heads were trained with a batch size of 112, an initial learning rate (LR) of 1e-4, the AdamW optimizer, random seed of 30, and cross-entropy loss.
+We use fivefold cross-validation for 50% shot fine-tuning.
 An early stopping criterion was applied, stopping training after 10 epochs if the validation loss shows no improvement.
-For the surgical tool presence detection tasks and the CholecT50 action recognition task, we used a linear head.
-For the surgical phase recognition tasks, we utilized a TCN [12] head and followed the two-stage training approach outlined in TeCNO [12].
+Surgical phase recognition To compare with other specialist models that are specifically tailored for this task, we adapt the video-level accuracy and Jaccard in the full fine-tuning setting.
+The accuracy for a video is computed by dividing the number of frames whose class has been correctly predicted by the total number of frames.
+The overall accuracy for a dataset is defined as the mean accuracy value over all videos, henceforth referred to as video-level accuracy [28, 29, 35, 27].
+The Jaccard is computed for each class and video, then averaged: first within videos, and secondly across classes [28, 29, 35, 27].
+During training, we utilized a TCN [12] head and followed the two-stage training approach outlined in TeCNO [12].
 In the first stage, we trained the backbone with a linear head to perform frame-wise classification without temporal context.
@@
 The checkpoint with the lowest validation loss was selected as the optimal checkpoint.
-For the SAR-RARP50 [48] action recognition task, we trained our model in two modes: online, where the SurgFM teacher model was used as backbone with a TCN head that could see only the information from previous frames; and offline, where the SurgFM teacher model was paired with an MS-TCN head [33] to incorporate both previous and future frame information for a fair comparison with the methods in [48].
-For the TCN head, we maintained the same model configuration and hyperparameters as in the surgical phase recognition tasks.
-For the MS-TCN head, we adopted a 3-stage, 11-layer configuration, following the basic model setup in [33].
-The model was trained using the Adam optimizer with an initial learning rate of 5e-4 for 70 epochs.
-Data scaling.
-In Sec.
-6 of the manuscript, we present the results of the data scaling experiments.
-For each scale (10%, 25%, 50%), we conducted a 5-fold cross-validation, with each fold randomly selected from the training dataset.
-The model configurations and hyperparameters were kept consistent across different data scales using the same settings as those used for the full training dataset.
-Appendix E Evaluation of SurgFM-Vid
-To evaluate the performance of our video classifier, SurgFM-Vid, we have split the Surg-3M dataset by videos, with 3369 videos for training and 825 videos for testing.
-Our vanilla SurgFM was trained on all the Surg-3M frames, therefore, to evaluate the video classification performance fairly, we have trained a new SurgFM model from scratch using only the frames of the 3369 videos in the training set that are accessible to all other classifiers in the comparison (i.e., 2.6 million frames instead of the usual 3.3 million frames in Surg-3M).
+Appendix E Evaluation of LemonFM-Vid
+To evaluate the performance of our video classifier, LemonFM-Vid, we have split the LEMON dataset by videos, with 3369 videos for training and 825 videos for testing.
+Our vanilla LemonFM was trained on all the LEMON frames, therefore, to evaluate the video classification performance fairly, we have trained a new LemonFM model from scratch using only the frames of the 3369 videos in the training set that are accessible to all other classifiers in the comparison.
 Appendix F Dataset and Licensing
 We adhere to the same practices as other datasets created from YouTube and various sources on the Web, such as ImageNet [13], Kinetics [9], YouTube-VIS [56], YouTube-8M [1], Insect-1M [42], Moments-in-Time [41], Tai-Chi-HD [49], HD-Villa-100M [55], and AVSpeech [14].
-Specifically, we provide the code to download and generate the Surg-3M dataset, a list of links to the original YouTube videos, and the corresponding annotations.
-Researchers working in academic institutions can request direct access to the full Surg-3M dataset in LMDB format for non-commercial purposes.
-We will also provide an online form for the YouTube video authors to opt out of our Surg-3M dataset.
-The Surg-3M dataset is provided under the Creative Commons Attribution 4.0 International (CC BY 4.0) license.
+Specifically, we provide the code to download and generate the LEMON dataset, a list of links to the original YouTube videos, and the corresponding annotations.
+Researchers working in academic institutions can request direct access to the LEMON dataset (1 fps version) in LMDB format for non-commercial purposes.
+We will also provide an online form for the YouTube video authors to opt out of our LEMON dataset.
+The LEMON dataset is provided under the Creative Commons Attribution 4.0 International (CC BY 4.0) license.